### v3

# Positive Semidefinite Supermartingales and Randomized Matrix Concentration Inequalities

Abstract
We present new concentration inequalities for either martingale dependent or exchangeable random symmetric matrices under a variety of tail conditions, encompassing now-standard Chernoff bounds to self-normalized heavy-tailed settings. These inequalities are often randomized in a way that renders them strictly tighter than existing deterministic results in the literature, are typically expressed in the Loewner order, and are sometimes valid at arbitrary data-dependent stopping times. Along the way, we explore the theory of positive semidefinite supermartingales and maximal inequalities, a natural matrix analog of scalar nonnegative supermartingales that is potentially of independent interest.
1 Introduction
Bounding the tail probability of the sum of independent and identically distributed (i.i.d.) real-valued random variables is one of the most crucial tasks in the nonasymptotic theory of statistics, for which a copious collection of methods has flourished [Boucheron et al., 2013]. This paper concerns concentration inequalities for sums of random symmetric matrices under martingale dependence or exchangeability (both generalizing the i.i.d. assumption).
Notable previous results include the study of Rademacher and random graph adjacency matrices by Oliveira [2009, 2010]; those in a series of seminal papers by Tropp [2012, 2015], who famously ports the techniques of cumulant generating functions (CGFs) and Chernoff-type bounds, which are central in scalar concentration, into the matrix regime; some new bounds by Mackey et al. [2014] using exchangeability; and a recent study on “anytime-valid” concentration inequalities by Howard et al. [2020, 2021] which improves and unifies several matrix Chernoff-type bounds. We return to several of these works later in the current paper.
To set the stage for our exposition, we quote a celebrated result by Mackey et al. [2014, Corollary 4.2] (and also Tropp [2012, Theorem 1.3]) which states that, if $d\times d$ random i.i.d. symmetric matrices $\mathbf{X}_{1},\dots,\mathbf{X}_{n}$ satisfy $(\mathbf{X}_{1}-\mathbb{E}\mathbf{X}_{1})^{2}\preceq\mathbf{B}$ (meaning that $\mathbf{B}-(\mathbf{X}_{1}-\mathbb{E}\mathbf{X}_{1})^{2}$ is positive semidefinite) almost surely, then
$\mathbb{P}\left(\lambda_{\max}\left(\frac{\mathbf{X}_{1}+\dots+\mathbf{X}_{n}}{n}-\mathbb{E}\mathbf{X}_{1}\right)\geqslant\sqrt{\frac{2\lambda_{\max}(\mathbf{B})\log(d/\alpha)}{n}}\right)\leqslant\alpha.$ (MHI)
Above, $\lambda_{\max}(\,\cdot\,)$ refers to the largest eigenvalue of a symmetric matrix, a typical scalarization of matrix-valued tails. When $d=1$, it recovers, with the same optimal constant, the scalar Hoeffding’s inequality as a special case; hence its name “matrix Hoeffding’s inequality”.
In this paper, we present numerous new techniques to strengthen this type of matrix concentration inequalities. Each of our new bounds enjoys a combination of some of the following virtues:
  1. 1.
    Anytime validity. We derive new inequalities that hold not only for a fixed number ($n$) of matrices but also a stopping time number of them [Howard et al., 2020].
  2. 2.
    Randomization. We losslessly tighten many of the previously known inequalities by randomly decreasing the tail threshold (decreasing the $a$ in (MHI) on the left hand side but not the right hand side), extending a technique that was recently explored by Ramdas and Manole [2023] for scalars.
  3. 3.
    Exchangeability. We go beyond the i.i.d. or martingale difference assumptions on random matrices and derive bounds that allow them to just be exchangeable instead.
  4. 4.
    Tighter (or no) scalarization. In some bounds we use the trace functional $\mathrm{tr}(\,\cdot\,)$ which is larger than the usual $\lambda_{\max}(\,\cdot\,)$ (e.g. in (MHI)), and introduce some bounds completely in the matrix order $\preceq$ without any scalarization, which is partly achieved by studying the maximal inequalities of matrix-valued sub- and super-martingales.
  5. 5.
    Heavy tails. Instead of assuming the matrices to be bounded or (sub)Gaussian, we sometimes allow them to only have a finite variance or $p$th moment, or be symmetric.
To showcase these virtues, let us pick two of the bounds we shall prove in the exposition that follows. Under the assumptions of the (MHI) above, our Theorem 5.5 states that
$\mathbb{P}\left(\lambda_{\max}\left(\frac{\gamma_{1}\mathbf{X}_{1}+\dots+\gamma_{\tau}\mathbf{X}_{\tau}}{\gamma_{1}+\dots+\gamma_{\tau}}-\mathbb{E}\mathbf{X}_{1}\right)\geqslant\frac{\log(dU/\alpha)+\frac{1}{2}\lambda_{\max}(\mathbf{B})\sum_{i=1}^{\tau}\gamma_{i}^{2}}{\gamma_{1}+\dots+\gamma_{\tau}}\right)\leqslant\alpha$ (1)
for any predictable positive tuning parameters $\{\gamma_{n}\}$, stopping time $\tau$, and an independent random variable $U\sim\text{Unif}_{(0,1)}$. The above equips (MHI) with virtues 1 and 2 without compromises, since replacing $U$ with 1 and taking $\tau=n$, $\gamma_{1}=\dots=\gamma_{n}=\sqrt{\frac{2\log(d/\alpha)}{n\lambda_{\max}(\mathbf{B})}}$ in above, one obtains the exact deviation bound $\sqrt{\frac{2\log(d/\alpha)\lambda_{\max}(\mathbf{B})}{n}}$ as in (MHI). A bound of very different flavor is that of our Theorem 4.17, which states that for exchangeable matrices $\mathbf{X}_{1},\dots,\mathbf{X}_{n}$ with variance $\mathbf{V}$,
$\mathbb{P}\left(\exists n\geqslant 1:\operatorname{abs}\left(\frac{\mathbf{X}_{1}+\dots+\mathbf{X}_{n}}{n}-\mathbb{E}\mathbf{X}_{1}\right)\npreceq\mathbf{A}\right)\leqslant\mathrm{tr}(\mathbf{V}\mathbf{A}^{-2})$ (2)
for any positive definite $\mathbf{A}$. This bound enjoys virtues 3, 4, and 5, sequentializing the matrix Chebyshev inequality $\mathbb{P}\left(\operatorname{abs}\left(\mathbf{X}_{1}-\mathbb{E}\mathbf{X}_{1}\right)\npreceq\mathbf{A}\right)\leqslant\mathrm{tr}(\mathbf{V}\mathbf{A}^{-2})$ by Ahlswede and Winter [2002, Appendix C] to an infinite time horizon.
Outline.
The rest of the paper is organized as follows. In Section 2, we briefly review the technical background (and challenges) of matrix concentration bounds as well as a collection of classical and recent scalar bounds that anticipate our matrix extensions. We present in Section 3 our first kind of bounds, derived via randomized matrix Markov, Chebyshev, and Chernoff inequalities. We develop and apply the theory of matrix-valued supermartingales and backward submartingales in Section 4, further obtaining bounds for i.i.d. and exchangeable matrices. Finally, in Section 5, we revisit scalar supermartingales (and the scalar processes they dominate) constructed via matrices which yield some other randomized bounds as well, concluding with a discussion on the respective advantages of matrix- and scalar- based sequential tests in both theoretical and empirical dimensions.
2 Preliminaries
2.1 Matrix Analysis
Let $\mathcal{S}_{d}$ denote the set of all $d\times d$ real-valued symmetric matrices, which is the only class of matrices considered in this paper. [^1] We only briefly discuss the case with non-symmetric and non-square matrices in Appendix D. These matrices are denoted by bold upper-case letters $\mathbf{A},\mathbf{B}$, etc; whereas $d$ dimensional (column) vectors by bold lower-case letters $\mathbf{u},\mathbf{v}$, etc. For $I\subseteq\mathbb{R}$, we denote by $\mathcal{S}_{d}^{I}$ the set of all real symmetric matrices whose eigenvalues are all in $I$. $\mathcal{S}_{d}^{[0,\infty)}$, the set of positive semidefinite and $\mathcal{S}_{d}^{(0,\infty)}$, the set of positive definite matrices are simply denoted by $\mathcal{S}_{d}^{+}$ and $\mathcal{S}_{d}^{++}$ respectively. The Loewner partial order is denoted $\preceq$, where $\mathbf{A}\preceq\mathbf{B}$ means $\mathbf{B}-\mathbf{A}$ is positive semidefinite, and $\mathbf{A}\prec\mathbf{B}$ means $\mathbf{B}-\mathbf{A}$ is positive definite.
Any function $f:I\to J$ (where $I,J\subseteq\mathbb{R}$) is identified canonically with a function $f:\mathcal{S}_{d}^{I}\to\mathcal{S}_{d}^{J}$ in the following way: for a diagonal matrix $\mathbf{\Sigma}\in\mathcal{S}_{d}^{I}$, $f(\mathbf{\Sigma})$ is obtained by applying $f$ entrywise; otherwise it is obtained by applying $f$ after orthogonal diagonalization, $f(\mathbf{U}^{\mathsf{T}}\mathbf{\Sigma}\mathbf{U})=\mathbf{U}^{\mathsf{T}}f(\mathbf{\Sigma})\mathbf{U}$. Notable matrix-to-matrix operators that recur in this paper defined this way include the matrix exponential $\exp:\mathcal{S}_{d}\to\mathcal{S}^{++}_{d}$, the matrix logarithm $\log:\mathcal{S}^{++}_{d}\to\mathcal{S}_{d}$, the matrix absolute value [^2] We use $\operatorname{abs}(\,\cdot\,)$ instead of ||$\cdot$|| ⋅ | to distinguish it from the determinant and other norms. $\operatorname{abs}:\mathcal{S}_{d}\to\mathcal{S}^{+}_{d}$, and powers $\mathbf{X}\mapsto\mathbf{X}^{k}$ for any $k\in\mathbb{R}$. Take heed that usually $\exp(\mathbf{A}+\mathbf{B})\neq\exp\mathbf{A}+\exp\mathbf{B}$ and $\log\mathbf{A}\mathbf{B}\neq\log\mathbf{A}+\log\mathbf{B}$, but a trace equality holds on the latter:
Fact 1 (Trace Logarithm Identity).
$\mathrm{tr}\log\mathbf{A}\mathbf{B}=\mathrm{tr}\log\mathbf{A}+\mathrm{tr}\log\mathbf{B}$.
[…]
It is rare to find matrix-to-matrix functions that are both monotone (i.e. preserving the $\preceq$ order) and convex (i.e. $f$ such that $f(\sum\lambda_{i}\mathbf{X}_{i})\preceq\sum{\lambda_{i}}f(\mathbf{X}_{i})$ if $\sum\lambda_{i}=1$ and each $\lambda_{i}\geqslant 0$; concavity is defined similarly). Scalar-to-scalar monotonicity ($I\to J$) does not imply matrix-to-matrix monotonicity ($\mathcal{S}_{d}^{I}\to\mathcal{S}_{d}^{J}$), and the same holds for convexity. For example, the matrix square $\mathbf{X}\mapsto\mathbf{X}^{2}$ is not monotone on the domain $\mathcal{S}_{d}^{+}$ and the matrix 4th power $\mathbf{X}\mapsto\mathbf{X}^{4}$ is neither monotone nor convex. The following are some of the few well-known nontrivial “operator-monotone” and “operator-convex” functions in the matrix analysis literature (cf. Bhatia [2013, Chapter V]).
Fact 2 (Operator-monotone and -concave Functions).
The matrix logarithm $\mathbf{X}\mapsto\log\mathbf{X}$ ($\mathcal{S}_{d}^{++}\to\mathcal{S}_{d}$) is monotone and concave. For $k\in(0,1)$, the matrix power $\mathbf{X}\mapsto\mathbf{X}^{k}$ ($\mathcal{S}_{d}^{+}\to\mathcal{S}_{d}^{+}$) is monotone and concave. For $k\in(1,2]$, the matrix power $\mathbf{X}\mapsto\mathbf{X}^{k}$ ($\mathcal{S}_{d}^{+}\to\mathcal{S}_{d}^{+}$ if $k\neq 2$; $\mathcal{S}_{d}\to\mathcal{S}_{d}^{+}$ if $k=2$) is convex.
[…]
Fact 3 (Trace Monotonicity).
Let $f:I\to\mathbb{R}$ be monotone increasing ($I\subseteq\mathbb{R}$). Then so is $\mathrm{tr}\circ f:\mathcal{S}_{d}^{I}\to\mathbb{R}$, i.e.,
$\mathbf{A}\preceq\mathbf{B}\implies\mathrm{tr}f(\mathbf{A})\leqslant\mathrm{tr}f(\mathbf{B}).$ (3)
Fact 4 (Trace Jensen).
For an interval $I\subseteq\mathbb{R}$, assume $f:I\to\mathbb{R}$ is convex. Then so is $\mathrm{tr}\circ f:\mathcal{S}_{d}^{I}\to\mathbb{R}$, i.e.,
$\mathrm{tr}f\left(\sum\lambda_{i}\mathbf{X}_{i}\right)\leqslant\mathrm{tr}\left(\sum\lambda_{i}f(\mathbf{X}_{i})\right),$ (4)
if $\sum\lambda_{i}=1$ and each $\lambda_{i}\geqslant 0$.
2.2 Concentration of Scalars
We briefly review some classical and recent scalar tail bounds in this subsection. Markov inequality, the foundation of a large family of scalar concentration inequalities, states that if $X$ is a nonnegative random variable,
$\mathbb{P}(X\geqslant a)\leqslant(\mathbb{E}X)a^{-1}$ (MI)
for any $a>0$. There have been many generalizations of (MI). Ville’s inequality [Ville, 1939] states that a similar tail bound holds uniformly over time for a nonnegative supermartingale $\{M_{n}\}$,
$\mathbb{P}(\exists n,\ M_{n}\geqslant a)\leqslant(\mathbb{E}M_{0})a^{-1},$ (VI)
or equivalently, $\mathbb{P}(M_{\tau}\geqslant a)\leqslant(\mathbb{E}M_{0})a^{-1}$ for any stopping time $\tau$ [Howard et al., 2020]. Ville’s inequality (VI) has been the workhorse for many recent time-uniform concentration inequalities (“confidence sequences”) [Howard et al., 2021, Waudby-Smith and Ramdas, 2023, 2020, Wang and Ramdas, 2023a, b] that have been favored by, for example, the multi-armed bandit literature.
Applying (MI) to the square of the sum of i.i.d. random variables $X_{1},\dots,X_{n}$ of mean $\mu$ and variance $\sigma^{2}$, we have the famous Chebyshev inequality
$\mathbb{P}\left(\left|\overline{X}_{n}-\mu\right|\geqslant a\right)\leqslant\frac{\sigma^{2}}{na^{2}},$ (CI)
where $\overline{X}_{n}$ is the arithmetic mean of $X_{1},\dots,X_{n}$. A perhaps less-known extension of (CI) is that if $X_{1},\dots,X_{n}$ have central $p$th moment $v_{p}$ ($1\leqslant p\leqslant 2$) instead of variance,
$\mathbb{P}\left(\left|\overline{X}_{n}-\mu\right|\geqslant a\right)\leqslant\frac{2^{2-p}v_{p}}{n^{p-1}a^{p}},$ (5)
which we prove in Section A.1 for completeness. In the other direction, if $X_{i}$’s are light-tailed and have moment generating function (MGF) $G(\gamma)=\mathbb{E}\mathrm{e}^{\gamma(X_{1}-\mu)}$, (MI) applied to the exponential of their mean yields the Chernoff-Hoeffding inequality
$\mathbb{P}(\overline{X}_{n}-\mu\geqslant a)\leqslant G(\gamma/n)^{n}\mathrm{e}^{-\gamma a}.$ (6)
Recently, Ramdas and Manole [2023] discovered some interesting strengthened forms of the bounds above. First, uniform randomization. It is found that if $U$ is a $\mathrm{Unif}_{(0,1)}$-distributed random variable independent from $X$ or $\{M_{n}\}$, (MI) and (VI) and be losslessly tightened into
$\mathbb{P}(X\geqslant Ua)\leqslant(\mathbb{E}X)a^{-1}$ (UMI)
and
$\mathbb{P}(M_{\tau}\geqslant Ua)\leqslant(\mathbb{E}M_{0})a^{-1}$ (UVI)
respectively. The downstream bounds including (CI), (5) and (6) can thus also be randomized into strictly tighter forms. Second, Ramdas and Manole [2023] introduce exchangeable generalizations of (CI) which partly shares the anytime virtue of (VI)-type bounds, stating that as long as the infinite sequence $X_{1},X_{2},\dots$ is exchangeable,
$\mathbb{P}\left(\exists n,\ \left|\overline{X}_{m}-\mu\right|\geqslant a\right)\leqslant\frac{\sigma^{2}}{a^{2}};$ (XCI)
and further, if i.i.d., (XCI) implies that for any fixed $N$,
$\mathbb{P}\left(\exists n\geqslant N,\ \left|\overline{X}_{m}-\mu\right|\geqslant a\right)\leqslant\frac{\sigma^{2}}{Na^{2}}.$ (7)
Extensions of (MI) and (VI) to random symmetric matrices are proved by Ahlswede and Winter [2002, Appendix C]; of (6) by Tropp [2012, 2015]. We will bring up these previous extensions again soon in the following. Our current paper is partly animated by the question if these novel bounds (UMI), (UVI), (XCI), and (7) by Ramdas and Manole [2023] have matrix siblings as well; and if yes, to what extent of generality. We shall show that there is indeed a plethora of matrix tail bounds that recover as $d=1$ special cases the results mentioned above, using classical matrix analysis tools plus some new probability techniques.
2.3 Templates of Our Bounds
We now sketch out the skeletons that our bounds may have, and discuss their relation. Let $\mathbf{X}_{1},\dots,\mathbf{X}_{n}$ be i.i.d. random symmetric matrices whose common distribution belongs to a family $\mathcal{P}$. Denote by $\mathbf{M}$ their common mean and $\overline{\mathbf{X}}_{n}$ their empirical average $\frac{1}{n}(\mathbf{X}_{1}+\dots+\mathbf{X}_{n})$. The following are the concentration events that we may prove to be of high probability. $\mathbf{A}\in\mathcal{S}^{+}_{d}$ denotes a constant matrix of deviation, and $r(\cdot)$ is a “randomization” map producing an $r(\mathbf{A})$ that is random but always smaller than $\mathbf{A}$ in Loewner order (the $Ua$ term in (UMI) is a one-dimensional example of $r(a)$). An arrow $A\to B$ denote the inclusion of events $A\subseteq B$, and hence also that $A$ is a stronger claim than $B$.
{tikzcd}
Most results in the past, for example (MHI), have been spectral bounds, which are weaker than many of the bounds we shall present in this paper.
3 Randomized Fixed-Time Matrix Tail Bounds
3.1 Randomized Matrix Markov Inequality
We begin our exposition with a matrix version of uniformly randomized Markov inequality (UMI) by Ramdas and Manole [2023]; a randomized version of the matrix Markov inequality by Ahlswede and Winter [2002, Theorem 12],
$\mathbb{P}(\mathbf{X}\npreceq\mathbf{A})\leqslant\mathrm{tr}((\mathbb{E}\mathbf{X})\mathbf{A}^{-1}),$ (MMI)
where $\mathbf{X}$ is an $\mathcal{S}_{d}^{+}$-valued random matrix and $\mathbf{A}\in\mathcal{S}_{d}^{++}$ is deterministic.
To achieve this, let us first define what it means for a random matrix to be uniformly and super-uniformly distributed. Note that it is easy to see that the randomizer $U$ in (UMI) and (UVI) can be replaced by any random variable that is stochastically larger; that is, one satisfying the following definition:
Definition 3.1.
We call a positive random scalar $U$ super-uniform if $\mathbb{P}(U\leqslant x)\leqslant x$ for all $x\geqslant 0$; that is, $U$ is stochastically greater than or equal to the uniform distribution $\operatorname{Unif}_{(0,1)}$.
This has two viable extensions to matrices, one in terms of the trace and the other of the spectral norm.
Definition 3.2.
We call an $\mathcal{S}_{d}^{+}$-valued random matrix $\mathbf{U}$ spectrally super-uniform if $\mathbb{P}(\mathbf{U}\nsucceq\mathbf{Y})\leqslant\lambda_{\max}(\mathbf{Y})$ for all $\mathbf{Y}\in\mathcal{S}_{d}^{+}$, and $\mathcal{Y}$-spectrally uniform if, further, equality holds for all $\mathbf{Y}\in\mathcal{Y}\subseteq\mathcal{S}_{d}^{+}$. We say $\mathbf{U}$ is trace super-uniform if $\mathbb{P}(\mathbf{U}\nsucceq\mathbf{Y})\leqslant\mathrm{tr}\mathbf{Y}$ for all $\mathbf{Y}\in\mathcal{S}_{d}^{+}$, and $\mathcal{Y}$-trace uniform if, further, equality holds for all $\mathbf{Y}\in\mathcal{Y}\subseteq\mathcal{S}_{d}^{+}$.
Note that every spectrally super-uniform matrix is a trace super-uniform matrix since $\lambda_{\max}(\mathbf{Y})\leqslant\mathrm{tr}\mathbf{Y}$. If $U$ is a super-uniform random scalar, then $U\mathbf{I}$ is a spectrally (hence trace) super-uniform random matrix, because $\mathbb{P}(U\mathbf{I}\nsucceq\mathbf{Y})=\mathbb{P}(U\leqslant\lambda_{\max}(\mathbf{Y}))\leqslant\lambda_{\max}(\mathbf{Y}).$
Straightforwardly the identity matrix $\mathbf{I}$ is spectrally (hence trace) super-uniform. If $U\sim\operatorname{Unif}_{(0,1)}$, $U\mathbf{I}$ is $\mathcal{S}_{d}^{[0,1]}$-spectrally uniform, and $\left\{\mathbf{v}\mathbf{v}^{\mathsf{T}}:\|\mathbf{v}\|\leqslant 1\right\}$-trace uniform. Having defined trace super-uniformity, we can now state a matrix extension of the uniformly randomized Markov’s inequality (UMI).
Theorem 3.3 (Uniformly Randomized Matrix Markov Inequality).
Let $\mathbf{X}$ be a random matrix taking values in $\mathcal{X}\subseteq\mathcal{S}_{d}^{+}$, and $\mathbf{U}$ a trace super-uniform random matrix taking values in $\mathcal{S}_{d}^{+}$ independent from $\mathbf{X}$. For any $\mathbf{A}\in\mathcal{S}_{d}^{++}$,
$\mathbb{P}(\mathbf{X}\npreceq\mathbf{A}^{1/2}\mathbf{U}\mathbf{A}^{1/2})\leqslant\mathrm{tr}((\mathbb{E}\mathbf{X})\mathbf{A}^{-1}).$ (UMMI)
The above holds with equality if and only if $\mathbf{U}$ is $\mathbf{A}^{-1/2}\mathcal{X}\mathbf{A}^{-1/2}$-trace uniform.
Proof.
By our assumptions,
$\mathbb{P}(\mathbf{X}\npreceq\mathbf{A}^{1/2}\mathbf{U}\mathbf{A}^{1/2})=\mathbb{P}(\mathbf{A}^{-1/2}\mathbf{X}\mathbf{A}^{-1/2}\npreceq\mathbf{U})=\mathbb{E}(\mathbb{P}(\mathbf{A}^{-1/2}\mathbf{X}\mathbf{A}^{-1/2}\npreceq\mathbf{U}|\mathbf{X}))$ (8)
[…]
concluding the proof. ∎
[…]
$\mathbb{P}(\lambda_{\max}(\mathbf{X})>\mathrm{tr}(\mathbf{A}\mathbf{U}))\leqslant\mathbb{P}(\mathrm{tr}\mathbf{X}>\mathrm{tr}(\mathbf{A}\mathbf{U}))\leqslant\mathbb{P}(\mathbf{X}\npreceq\mathbf{A}^{1/2}\mathbf{U}\mathbf{A}^{1/2}),$ (10)
subject to the same upper bound $\mathrm{tr}((\mathbb{E}\mathbf{X})\mathbf{A}^{-1})$.
The following is an example where the inequality of (UMMI) holds with equality. We take $\mathbf{U}=U\mathbf{I}$ where $U\sim\operatorname{Unif}_{(0,1)}$, and $\mathbf{A}\in\mathcal{S}_{d}^{++}$. Let $\mathbf{x}$ be a random vector in the ellipsoid $\{\mathbf{x}:\mathbf{x}^{\mathsf{T}}\mathbf{A}^{-1}\mathbf{x}\leqslant 1\}\subseteq\mathbb{R}^{d}$. Then,
$\mathbb{P}(\|\mathbf{A}^{-1/2}\mathbf{x}\|^{2}\leqslant U)=\mathbb{P}(\mathbf{x}\mathbf{x}^{\mathsf{T}}\npreceq U\mathbf{A})=\mathrm{tr}(\mathbf{A}^{-1}\mathbb{E}(\mathbf{x}\mathbf{x}^{\mathsf{T}}))=\mathbb{E}\|\mathbf{A}^{-1/2}\mathbf{x}\|^{2},$ (11)
where the second equality above is (UMMI) applied on the random matrix $\mathbf{x}\mathbf{x}^{\mathsf{T}}$. It is not hard to see here the condition for equality is met. Intuitively, this hints that (UMMI) is tighter when the random matrix $\mathbf{X}$ is approximately of lower rank, a point that we shall return to later in this paper with statistical examples.
Concluding this section, we remark on trace super-uniformity that clearly $U\mathbf{I}+\mathbf{Y}$ is trace super-uniform when $U\sim\operatorname{Unif}_{(0,1)}$ and $\mathbf{Y}\succeq 0$, and this provides an interesting class for our purposes. It may be interesting to fully characterize the set of all trace super-uniform matrices, but we leave it as future work.
3.2 Randomized Matrix Chebyshev Inequalities
Recall that the variance of a random symmetric matrix $\mathbf{X}$ is defined simply as $\mathbf{Var}(\mathbf{X})=\mathbb{E}(\mathbf{X}-\mathbb{E}\mathbf{X})^{2}$. Applying Theorem 3.3 to the square of one square-integrable matrix and the square of the sum of several i.i.d. square-integrable matrices, we have the following randomization of the one-matrix Chebyshev inequality by Ahlswede and Winter [2002, Theorem 14],
$\mathbb{P}(\operatorname{abs}(\mathbf{X}-\mathbb{E}\mathbf{X})\npreceq\mathbf{A})\leqslant\mathrm{tr}(\mathbf{Var}(\mathbf{X})\mathbf{A}^{-2})$ (12)
and the implied law of large numbers for the average $\overline{\mathbf{X}}_{n}$ of $n$ i.i.d. symmetric matrices with mean $\mathbf{M}$ and variance $\mathbf{V}$ by Ahlswede and Winter [2002, Theorem 15],
$\mathbb{P}(\operatorname{abs}(\overline{\mathbf{X}}_{n}-\mathbf{M})\npreceq\mathbf{A})\leqslant n^{-1}\mathrm{tr}(\mathbf{V}\mathbf{A}^{-2}).$ (MCI)
Lemma 3.4 (Uniformly Randomized One-Matrix Chebyshev Inequality).
Let $\mathbf{X}$ be a random matrix taking values in $\mathcal{S}_{d}$, $\mathbf{U}$ a trace super-uniform random matrix taking values in $\mathcal{S}_{d}^{+}$ independent from $\mathbf{X}$. For any $\mathbf{A}\in\mathcal{S}_{d}^{++}$,
$\mathbb{P}(\operatorname{abs}(\mathbf{X}-\mathbb{E}\mathbf{X})\npreceq(\mathbf{A}\mathbf{U}\mathbf{A})^{1/2})\leqslant\mathrm{tr}(\mathbf{Var}(\mathbf{X})\mathbf{A}^{-2}).$ (13)
Proof.
By Fact 2, $\mathbb{P}(\operatorname{abs}(\mathbf{X}-\mathbb{E}\mathbf{X})\npreceq(\mathbf{A}\mathbf{U}\mathbf{A})^{1/2})\leqslant\mathbb{P}((\mathbf{X}-\mathbb{E}\mathbf{X})^{2}\npreceq\mathbf{A}\mathbf{U}\mathbf{A})$. By Theorem 3.3, we have $\mathbb{P}((\mathbf{X}-\mathbb{E}\mathbf{X})^{2}\npreceq\mathbf{A}\mathbf{U}\mathbf{A})\leqslant\mathrm{tr}(\mathbf{Var}(\mathbf{X})\mathbf{A}^{-2})$. ∎
We can immediately apply Lemma 3.4 to multiple matrices, obtaining the following weak-law-of-large-numbers style inequality that shrinks when $n$ gets larger. Recall that the anticommutator between two sqaure matrices $\mathbf{A},\mathbf{B}$ is defined as $\{\mathbf{A},\mathbf{B}\}=\mathbf{A}\mathbf{B}+\mathbf{B}\mathbf{A}$.
Theorem 3.5 (Uniformly Randomized $n$-Matrix Chebyshev Inequality).
Let $\mathbf{X}_{1},\dots,\mathbf{X}_{n}$ be random matrices taking values in $\mathcal{S}_{d}$ with common mean matrix $\mathbb{E}\mathbf{X}_{i}=\mathbf{M}$ and variance matrix upper bound $\mathbf{Var}(\mathbf{X}_{i})\preceq\mathbf{V}$, such that $\mathbb{E}\{\mathbf{X}_{i}-\mathbf{M},\mathbf{X}_{j}-\mathbf{M}\}=0$ for any $i\neq j$; $\mathbf{U}$ a trace super-uniform random matrix taking values in $\mathcal{S}_{d}^{+}$ independent from them all. Then, for $\overline{\mathbf{X}}_{n}=\frac{1}{n}(\mathbf{X}_{1}+\dots+\mathbf{X}_{n})$, we have
$\mathbf{Var}(\overline{\mathbf{X}}_{n})\preceq n^{-1}\mathbf{V}.$ (14)
Consequently, for any $\mathbf{A}\in\mathcal{S}_{d}^{++}$.
$\mathbb{P}(\operatorname{abs}(\overline{\mathbf{X}}_{n}-\mathbf{M})\npreceq(\mathbf{A}\mathbf{U}\mathbf{A})^{1/2})\leqslant n^{-1}\mathrm{tr}(\mathbf{V}\mathbf{A}^{-2}),$ (UMCI)
Proof.
Note that
$\small\mathbf{Var}(\overline{\mathbf{X}}_{n})=n^{-2}\mathbb{E}\left(\sum_{i=1}^{n}(\mathbf{X}_{i}-\mathbf{M})\right)^{2}=n^{-2}\left(\sum_{i=1}^{n}\mathbf{Var}(\mathbf{X}_{i})+\sum_{i<j}\mathbb{E}\{\mathbf{X}_{i}-\mathbf{M},\mathbf{X}_{j}-\mathbf{M}\}\right)\preceq n^{-1}\mathbf{V},$ (15)
and apply Lemma 3.4 to conclude the proof. ∎
It is easy to see that (UMCI) recovers as special cases (MCI) and (CI).
Remark 3.6.
The assumption $\mathbb{E}\{\mathbf{X}_{i}-\mathbf{M},\mathbf{X}_{j}-\mathbf{M}\}=0$ for $i\neq j$ is satisfied when $\mathbf{X}_{1},\mathbf{X}_{2},\dots$ are pairwise uncorrelated, which is in turn satisfied under martingale difference assumption (i.e., $\{\mathbf{X}_{i}\}$ is adapted to some filtration $\{\mathcal{F}_{i}\}$ with $\mathbb{E}(\mathbf{X}_{i}|\mathcal{F}_{i-1})=\mathbf{M}$ and $\mathbf{Var}(\mathbf{X}_{i}|\mathcal{F}_{i-1})\preceq\mathbf{V}$), which is in turn implied if they are i.i.d.
For $p\in[1,2)$, a $p$th moment bound can lead to an infinite variance version of (13), whose non-randomized ($\mathbf{U}=\mathbf{I}$) version is mentioned by Tu [2017].
Lemma 3.7 (Uniformly Randomized One-Matrix $p$-Chebyshev Inequality).
Let $\mathbf{X}$ be a random matrix taking values in $\mathcal{S}_{d}$ that has a finite $p$th central moment $\mathbf{V}_{p}:=\mathbb{E}(\operatorname{abs}(\mathbf{X}-\mathbb{E}\mathbf{X}))^{p}$ and $\mathbf{U}$ a trace super-uniform random matrix taking values in $\mathcal{S}_{d}^{+}$ independent from $\mathbf{X}$. For any $\mathbf{A}\in\mathcal{S}_{d}^{++}$,
$\mathbb{P}(\operatorname{abs}(\mathbf{X}-\mathbb{E}\mathbf{X})\npreceq(\mathbf{A}^{p/2}\mathbf{U}\mathbf{A}^{p/2})^{1/p})\leqslant\mathrm{tr}(\mathbf{V}_{p}\mathbf{A}^{-p}).$ (16)
Proof.
By Fact 2, $\mathbb{P}(\operatorname{abs}(\mathbf{X}-\mathbb{E}\mathbf{X})\npreceq(\mathbf{A}^{p/2}\mathbf{U}\mathbf{A}^{p/2})^{1/p})\leqslant\mathbb{P}((\operatorname{abs}(\mathbf{X}-\mathbb{E}\mathbf{X}))^{p}\npreceq\mathbf{A}^{p/2}\mathbf{U}\mathbf{A}^{p/2})$. By Theorem 3.3, we have $\mathbb{P}((\operatorname{abs}(\mathbf{X}-\mathbb{E}\mathbf{X}))^{p}\npreceq\mathbf{A}^{p/2}\mathbf{U}\mathbf{A}^{p/2})\leqslant\mathrm{tr}(\mathbf{V}_{p}\mathbf{A}^{-p})$. ∎
However, an $L^{p}$ analogy of (UMCI) (i.e. a matrix version of the $p$-Chebyshev inequality (5)) is not straightforward. This is largely because while the contraction of variance rule (14) exists for matrices, we do not know of any matrix version of $p$th moment contraction rule. In fact, the scalar $p$th moment contraction rule was only recently established (by Cherapanamjeri et al. [2022, Lemma 4.2] and Wang et al. [2021, Lemma 7], see the proof of (5) in Section A.1). Section A.2 contains a conjectured result along these lines, as well as a looser result that is the best we can currently achieve.
3.3 Randomized Matrix Chernoff Bounds
We now establish the classical Chernoff-type bounds with randomization for light-tailed random matrices that have finite moment generating functions (MGFs).
Lemma 3.8 (Uniformly Randomized One-Matrix Chernoff Bound).
Let $\mathbf{X}$ be a random matrix taking values in $\mathcal{S}_{d}$, $\mathbf{U}$ a trace super-uniform random matrix taking values in $\mathcal{S}_{d}^{+}$ independent from $\mathbf{X}$. For any $\mathbf{A}\in\mathcal{S}_{d}$ and real number $\gamma$,
$\mathbb{P}\left(\mathbf{X}\npreceq\frac{1}{2\gamma}\log(\mathrm{e}^{\gamma\mathbf{A}}\mathbf{U}\mathrm{e}^{\gamma\mathbf{A}})\right)\leqslant\mathrm{tr}(\mathrm{e}^{-2\gamma\mathbf{A}}\,\mathbb{E}\mathrm{e}^{2\gamma\mathbf{X}}).$ (17)
Proof.
By Fact 2, we have $\mathbb{P}\left(2\gamma\mathbf{X}\npreceq\log(\mathrm{e}^{\gamma\mathbf{A}}\mathbf{U}\mathrm{e}^{\gamma\mathbf{A}})\right)\leqslant\mathbb{P}(\mathrm{e}^{2\gamma\mathbf{X}}\npreceq\mathrm{e}^{\gamma\mathbf{A}}\mathbf{U}\mathrm{e}^{\gamma\mathbf{A}})$. By Theorem 3.3, we have $\mathbb{P}(\mathrm{e}^{2\gamma\mathbf{X}}\npreceq\mathrm{e}^{\gamma\mathbf{A}}\mathbf{U}\mathrm{e}^{\gamma\mathbf{A}})\leqslant\mathrm{tr}(\mathrm{e}^{-2\gamma\mathbf{A}}\,\mathbb{E}\mathrm{e}^{2\gamma\mathbf{X}})$, concluding the proof. ∎
The seemingly indigestible tail threshold $\frac{1}{2\gamma}\log(\mathrm{e}^{\gamma\mathbf{A}}\mathbf{U}\mathrm{e}^{\gamma\mathbf{A}})$ in (17) can be largely simplified in any of the following scenarios. First, one can simply take the trace on the tail event above, obtaining due to Fact 1 and the monotonicity of the trace,
$\mathbb{P}\left(\mathrm{tr}\mathbf{X}>\mathrm{tr}\mathbf{A}+\frac{1}{2\gamma}\mathrm{tr}\log\mathbf{U}\right)\leqslant\mathbb{P}\left(\mathbf{X}\npreceq\frac{1}{2\gamma}\log(\mathrm{e}^{\gamma\mathbf{A}}\mathbf{U}\mathrm{e}^{\gamma\mathbf{A}})\right),$ (18)
without any additional assumptions; Second, if $\mathbf{U}=U\mathbf{I}$ where $U$ is scalar super-uniform, $\frac{1}{2\gamma}\log(\mathrm{e}^{\gamma\mathbf{A}}\mathbf{U}\mathrm{e}^{\gamma\mathbf{A}})=\mathbf{A}+\frac{1}{2\gamma}\mathbf{I}\log U$; Third, if $\mathbf{A}=a\mathbf{I}$, $\frac{1}{2\gamma}\log(\mathrm{e}^{\gamma\mathbf{A}}\mathbf{U}\mathrm{e}^{\gamma\mathbf{A}})=a\mathbf{I}+\frac{1}{2\gamma}\log\mathbf{U}$.
It is known that, unlike the scalar case, the one-matrix MGF is not linearly additive. Once again, the complication is partially resolved when the trace is applied.
Fact 5 (Lemma 3.4 of Tropp [2012]).
Let $\mathbf{Z}_{1},\dots,\mathbf{Z}_{n}$ be independent random matrices taking values in $\mathcal{S}_{d}$. Then,
$\mathrm{tr}\left\{\mathbb{E}\exp\left(\sum\mathbf{Z}_{i}\right)\right\}\leqslant\mathrm{tr}\exp\left(\sum\log\mathbb{E}\mathrm{e}^{\mathbf{Z}_{i}}\right).$ (19)
This implies the following bound that recovers as special cases both the scalar Chernoff-Hoeffding inequality (6) when taking $d=\mathbf{U}=1$, and the “master tail bound” on $\lambda_{\max}(\overline{\mathbf{X}}_{n}-\mathbf{M})$ by Tropp [2012, Theorem 3.6] when taking $\mathbf{U}=\mathbf{I}$.
Theorem 3.9 (Uniformly Randomized Matrix Chernoff-Hoeffding Inequality).
Let $\mathbf{X}_{1},\dots,\mathbf{X}_{n}$ be i.i.d. random matrices taking values in $\mathcal{S}_{d}$ with common mean matrix $\mathbb{E}\mathbf{X}_{1}=\mathbf{M}$ and matrix-valued MGF $\mathbf{G}(\gamma)=\mathbb{E}\mathrm{e}^{\gamma(\mathbf{X}_{1}-\mathbf{M})}$; $\mathbf{U}$ a trace super-uniform random matrix taking values in $\mathcal{S}_{d}^{+}$ independent from them all. Let $\overline{\mathbf{X}}_{n}=\frac{1}{n}(\mathbf{X}_{1}+\dots+\mathbf{X}_{n})$. For any $a,t>0$,
$\mathbb{P}\left(\overline{\mathbf{X}}_{n}-\mathbf{M}\npreceq a\mathbf{I}+\frac{\log\mathbf{U}}{\gamma}\right)\leqslant\mathrm{tr}(\mathrm{e}^{n\log\mathbf{G}(\gamma/n)})\mathrm{e}^{-\gamma a}.$ (20)
Proof.
By Fact 5, $\mathrm{tr}(\mathbb{E}\mathrm{e}^{\gamma(\overline{\mathbf{X}}_{n}-\mathbf{M})})\leqslant\mathrm{tr}(\mathrm{e}^{n\log\mathbf{G}(\gamma/n)})$. Using Lemma 3.8 on $\gamma(\overline{\mathbf{X}}_{n}-\mathbf{M})$ with $\mathbf{A}=a\mathbf{I}$ concludes the proof. ∎
Again, the bound remains valid if only an upper bound of the matrix MGF $\mathbb{E}\mathrm{e}^{\gamma(\mathbf{X}_{1}-\mathbf{M})}\preceq\mathbf{G}(\gamma)$ or (more commonly) the weaker log-MGF bound $\log\mathbb{E}\mathrm{e}^{\gamma(\mathbf{X}_{1}-\mathbf{M})}\preceq\log\mathbf{G}(\gamma)$ is known. Table 1 contains some important known matrix (log-)MGFs (see Tropp [2012, 2015]). An asymmetric Hoeffding bound is missing here but shall be discussed in Section 5 later since it requires some further techniques beyond the matrix log-MGF.
Name Condition $\mathbf{G}(\gamma)$ $\mathrm{tr}(\mathrm{e}^{n\log\mathbf{G}(\gamma/n)})$
Rademacher $\mathbf{X}_{1}-\mathbf{M}=R_{1}\mathbf{C},R_{1}\sim\operatorname{rad}_{1/2}$ $\preceq\mathrm{e}^{\gamma^{2}\mathbf{C}^{2}/2}$ $\leqslant\mathrm{tr}(\mathrm{e}^{\gamma^{2}\mathbf{C}^{2}/2n})$
Uni-Gaussian $\mathbf{X}_{1}-\mathbf{M}=G_{1}\mathbf{C},G_{1}\sim\mathcal{N}_{0,1}$ $=\mathrm{e}^{\gamma^{2}\mathbf{C}^{2}/2}$ $=\mathrm{tr}(\mathrm{e}^{\gamma^{2}\mathbf{C}^{2}/2n})$
Bennett I $\lambda_{\max}(\mathbf{X}_{1}-\mathbf{M})\leqslant 1,\mathbf{Var}(\mathbf{X}_{1})=\mathbf{V}$ $\preceq\mathrm{e}^{(\mathrm{e}^{\gamma}-\gamma-1)\mathbf{V}}$ $\leqslant\mathrm{tr}(\mathrm{e}^{n(\mathrm{e}^{\gamma/n}-\gamma/n-1)\mathbf{V}})$
Bennett II $\lambda_{\max}(\mathbf{X}_{1}-\mathbf{M})\leqslant 1,\mathbf{Var}(\mathbf{X}_{1})\preceq\mathbf{V}$ $\leqslant\mathrm{tr}(\mathrm{e}^{n(\mathrm{e}^{\gamma/n}-\gamma/n-1)\mathbf{V}})$
Symmetric Hoeffding $(\mathbf{X}_{1}-\mathbf{M})^{2}\preceq\mathbf{B},\ \mathbf{X}_{1}-\mathbf{M}\stackrel{{\scriptstyle d}}{{=}}\mathbf{M}-\mathbf{X}_{1}$ $\leqslant\mathrm{tr}(\mathrm{e}^{\gamma^{2}\mathbf{B}/2n})$
Table 1: Some important matrix MGFs and log-MGFs. Here $\mathbf{B},\mathbf{C},\mathbf{V}$ are deterministic matrices, and =$\stackrel{{\scriptstyle d}}{{=}}$ means two random matrices having the same distribution. The proofs are in Section B.1.
4 Maximal Inequalities of Matrix (Super-/Sub-)Martingales
Some interesting tail bounds are obtained if we consider matrix-valued supermartingales and submartingales and their maximal inequalities. These processes can be defined on forward or backward filtrations and in both these cases, well-known scalar-valued maximal inequalities can be extended to matrices.
4.1 Matrix-Valued Supermartingales
Let $\{\mathbf{Y}_{n}\}$ be an $\mathcal{S}_{d}$-valued stochastic process adapted to the filtration $\{\mathcal{F}_{n}\}$. [^3] We shall often omit the filtration $\{\mathcal{F}_{n}\}$ when unambiguous from the context. We say that $\{\mathbf{Y}_{n}\}$ is a supermartingale if $\mathbb{E}(\mathbf{Y}_{n}|\mathcal{F}_{n-1})\preceq\mathbf{Y}_{n-1}$ holds for every $n$. Before we proceed, note that the order $\preceq$ is preserved by addition, limit, and Lebesgue integration (hence including convergent infinite sums, expectations, conditional expectations, etc., essentially because $\mathcal{S}^{+}=\{\mathbf{A}:\mathbf{A}\succeq 0\}$ is closed both topologically and under $+$). See also a “dominated integrability” lemma we prove in Section B.2.
Matrix supermartingales are related to scalar supermartingales in the following ways. First, the monotonicity of trace implies that if $\{\mathbf{Y}_{n}\}$ is an $\mathcal{S}_{d}$-valued supermartingale, its trace process $\{\mathrm{tr}(\mathbf{Y}_{n})\}$ is a real-valued supermartingale. Second, we have the following lemma.
Lemma 4.1.
An $\mathcal{S}_{d}$-valued process $\{\mathbf{Y}_{n}\}$ is a supermartingale if and only if for every deterministic vector $\mathbf{v}\in\mathbb{R}^{d}$, the scalar-valued process $\{\mathbf{v}^{\mathsf{T}}\mathbf{Y}_{n}\mathbf{v}\}$ is a supermartingale.
Proof.
$\{\mathbf{Y}_{n}\}$ is a supermartingale (21)
$\iff$ For all $n$, $\mathbf{Y}_{n-1}-\mathbb{E}(\mathbf{Y}_{n}|\mathcal{F}_{n-1})\succeq 0$ (22)
$\iff$ For all $n$ and $\mathbf{v}$, $\mathbf{v}^{\mathsf{T}}(\mathbf{Y}_{n-1}-\mathbb{E}(\mathbf{Y}_{n}|\mathcal{F}_{n-1}))\mathbf{v}\geqslant 0$ (23)
$\iff$ For all $n$ and $\mathbf{v}$, $\mathbf{v}^{\mathsf{T}}\mathbf{Y}_{n-1}\mathbf{v}-\mathbb{E}(\mathbf{v}^{\mathsf{T}}\mathbf{Y}_{n}\mathbf{v}|\mathcal{F}_{n-1})\geqslant 0$ (24)
$\iff$ For all $\mathbf{v}$, $\{\mathbf{v}^{\mathsf{T}}\mathbf{Y}_{n}\mathbf{v}\}$ is a supermartingale. (25)
This completes the proof. ∎
This leads to an analog of Doob’s martingale convergence theorem for positive semidefinite matrix supermartingales.
Theorem 4.2 (Matrix Supermartingale Convergence Theorem).
Any $\mathcal{S}_{d}^{+}$-valued supermartingale $\{\mathbf{Y}_{n}\}$ converges almost surely to an $\mathcal{S}_{d}^{+}$-valued random matrix $\mathbf{Y}_{\infty}$ and $\mathbb{E}\mathbf{Y}_{\infty}\preceq\mathbb{E}\mathbf{Y}_{0}$.
Proof.
For any $\mathbf{v}\in\mathbb{R}^{d}$, the process $\{\mathbf{v}^{\mathsf{T}}\mathbf{Y}_{n}\mathbf{v}\}$ is a scalar-valued nonnegative supermartingale, which converges a.s. to an integrable random scalar due to Doob’s martingale convergence theorem (see e.g. Corollary 11.5 in Klenke [2013]). First, setting $\mathbf{v}$ to $(0,\dots,0,1,0,\dots,0)^{\mathsf{T}}$ we see that $\mathbf{Y}_{n}$’s diagonal converges to integrable scalars. Then, setting $\mathbf{v}$ to $(0,\dots,0,1,0,\dots,0,1,0,\dots,0)^{\mathsf{T}}$ we see that the rest of $\mathbf{Y}_{n}$’s entries also converge to integrable scalars. The limit $\mathbf{Y}_{\infty}=\lim_{n\to\infty}\mathbf{Y}_{n}$ is in $\mathcal{S}_{d}^{+}$ because $\mathcal{S}_{d}^{+}$ is a closed set. Finally, for any $\mathbf{v}\in\mathbb{R}^{d}$,
$\mathbf{v}^{\mathsf{T}}(\mathbb{E}\mathbf{Y}_{\infty})\mathbf{v}=\lim_{n\to\infty}\mathbb{E}(\mathbf{v}^{\mathsf{T}}\mathbf{Y}_{n}\mathbf{v})\leqslant\mathbb{E}(\mathbf{v}^{\mathsf{T}}\mathbf{Y}_{0}\mathbf{v})=\mathbf{v}^{\mathsf{T}}(\mathbb{E}\mathbf{Y}_{0})\mathbf{v},$ (26)
so $\mathbb{E}\mathbf{Y}_{\infty}\preceq\mathbb{E}\mathbf{Y}_{0}$. ∎
Next, we see that using a similar trick, any scalar version of the optional stopping (sampling) theorems extends to symmetric matrices as well. For example:
Lemma 4.3 (Matrix Optional Stopping).
Let $\{\mathbf{Y}_{n}\}$ be an $\mathcal{S}_{d}$-valued supermartingale, and $\tau$ be a stopping time on $\{\mathcal{F}_{n}\}$. Then,
  1. 1.
    $\{\mathbf{Y}_{n\wedge\tau}\}$ is also a supermartingale on $\{\mathcal{F}_{n}\}$.
  2. 2.
    If $\tau\leqslant N$ almost surely for some $N$, then $\mathbb{E}\mathbf{Y}_{\tau}\preceq\mathbb{E}\mathbf{Y}_{0}$.
  3. 3.
    If $\{\mathbf{Y}_{n}\}$ is $\mathcal{S}_{d}^{+}$-valued, then $\mathbb{E}\mathbf{Y}_{\tau}\preceq\mathbb{E}\mathbf{Y}_{0}$.
Proof.
1. For any $\mathbf{v}\in\mathbb{R}^{d}$, the scalar process $\{\mathbf{v}^{\mathsf{T}}\mathbf{Y}_{n}\mathbf{v}\}$ is a supermartingale on $\{\mathcal{F}_{n}\}$ and therefore so is the stopped process $\{\mathbf{v}^{\mathsf{T}}\mathbf{Y}_{n\wedge\tau}\mathbf{v}\}$ due to the optional stopping theorem (e.g. Theorem 10.15 of Klenke [2013]). This implies that $\{\mathbf{Y}_{n\wedge\tau}\}$ is a matrix supermartingale on $\{\mathcal{F}_{n}\}$. 2. For any $\mathbf{v}\in\mathbb{R}^{d}$, we apply the optional sampling theorem (e.g. Theorem 10.11(i) of Klenke [2013]) to the scalar-valued supermartingale $\{\mathbf{v}^{\mathsf{T}}\mathbf{Y}_{n}\mathbf{v}\}$ to obtain that $\mathbf{v}^{\mathsf{T}}(\mathbb{E}\mathbf{Y}_{\tau})\mathbf{v}=\mathbb{E}(\mathbf{v}^{\mathsf{T}}\mathbf{Y}_{\tau}\mathbf{v})\leqslant\mathbb{E}(\mathbf{v}^{\mathsf{T}}\mathbf{Y}_{0}\mathbf{v})=\mathbf{v}^{\mathsf{T}}(\mathbb{E}\mathbf{Y}_{0})\mathbf{v}$. Therefore $\mathbb{E}\mathbf{Y}_{\tau}\preceq\mathbb{E}\mathbf{Y}_{0}$. 3. Apply Theorem 4.2 to the $\mathcal{S}_{d}^{+}$-valued stopped supermartingale $\{\mathbf{Y}_{\tau\wedge n}\}$, we have $\mathbb{E}\mathbf{Y}_{\tau\wedge\infty}\preceq\mathbb{E}\mathbf{Y}_{\tau\wedge 0}$, i.e., $\mathbb{E}\mathbf{Y}_{\tau}\preceq\mathbb{E}\mathbf{Y}_{0}$. ∎
This gives rise to the matrix versions of (VI) and (UVI) via some manipulation of crossing events.
Theorem 4.4 ((Uniformly Randomized) Matrix Ville’s Inequality).
Let $\{\mathbf{Y}_{n}\}$ be an $\mathcal{S}_{d}^{+}$-valued supermartingale on $\{\mathcal{F}_{n}\}$, and $\tau$ be a stopping time on $\{\mathcal{F}_{n}\}$. Let $\mathbf{U}$ be a trace super-uniform random matrix taking values in $\mathcal{S}_{d}^{+}$ independent from $\mathcal{F}_{\infty}$. Then, for any $\mathbf{A}\in\mathcal{S}_{d}^{++}$,
$\mathbb{P}(\mathbf{Y}_{\tau}\npreceq\mathbf{A}^{1/2}\mathbf{U}\mathbf{A}^{1/2})\leqslant\mathrm{tr}((\mathbb{E}\mathbf{Y}_{0})\mathbf{A}^{-1}).$ (UMVI)
Consequently,
$\mathbb{P}(\exists n,\;\mathbf{Y}_{n}\npreceq\mathbf{A})\leqslant\mathrm{tr}((\mathbb{E}\mathbf{Y}_{0})\mathbf{A}^{-1}).$ (MVI)
Proof.
First, $\mathbb{E}\mathbf{Y}_{\tau}\preceq\mathbb{E}\mathbf{Y}_{0}$ due to Part 3 of Lemma 4.3. By Theorem 3.3 and the monotinicity of the trace,
$\mathbb{P}(\mathbf{Y}_{\tau}\npreceq\mathbf{A}^{1/2}\mathbf{U}\mathbf{A}^{1/2})\leqslant\mathrm{tr}(\mathbf{A}^{-1/2}(\mathbb{E}\mathbf{Y}_{\tau})\mathbf{A}^{-1/2})\leqslant\mathrm{tr}(\mathbf{A}^{-1/2}(\mathbb{E}\mathbf{Y}_{0})\mathbf{A}^{-1/2}),$ (27)
which is (UMVI). Next, we define the stopping time $\nu:=\inf\{n:\mathbf{Y}_{n}\npreceq\mathbf{A}\}$. Setting $\mathbf{U}$ to $\mathbf{I}$ and $\tau$ to $\nu\wedge N$ in (UMVI), we have
$\mathbb{P}(\nu\leqslant N)=\mathbb{P}(\mathbf{Y}_{\nu\wedge N}\npreceq\mathbf{A})\leqslant\mathrm{tr}((\mathbb{E}\mathbf{Y}_{0})\mathbf{A}^{-1})$ (28)
for any $N\in\mathbb{N}$. Letting $N\to\infty$, we have
$\mathbb{P}(\nu<\infty)\leqslant\mathrm{tr}((\mathbb{E}\mathbf{Y}_{0})\mathbf{A}^{-1})$ (29)
which concludes the proof of (MVI). ∎
We remark that while there are two equivalent ways of stating the Ville’s inequality, the “any-time” statement (VI) and the stopping time statement (stated after (VI)), it appears that randomization only applies to the latter [Ramdas and Manole, 2023]. The same applies to our matrix versions of Ville’s inequality. Finally, we remark that some other maximal inequalities for scalar sub- and super-martingales can be extended to matrices as well, of which we derive a few in Appendix C.
4.2 Matrix Generalizations of Classical Test Supermartingales
Recall that scalar-valued nonnegative supermartingales are often constructed incrementally by multiplying each time a nonnegative variable with conditional expectation $\leqslant 1$. Analogously, we may use the following straightforward fact to construct $\mathcal{S}_{d}^{+}$-valued supermartingales.
Lemma 4.5.
Let $\mathbf{E}_{n}$ be $\mathcal{S}_{d}^{+}$-valued and $\mathcal{F}_{n}$-measurable, $\mathbf{F}_{n}$ be $\mathcal{S}_{d}^{++}$-valued and $\mathcal{F}_{n-1}$-measurable, such that $\mathbb{E}(\mathbf{E}_{n}|\mathcal{F}_{n-1})\preceq\mathbf{F}_{n}^{-1}$ almost surely. Then, the following $\mathcal{S}_{d}^{+}$-valued process is a supermartingale on $\{\mathcal{F}_{n}\}$:
$\mathbf{Y}_{n}=\sqrt{\mathbf{F}_{1}}\sqrt{\mathbf{E}_{1}}\sqrt{\mathbf{F}_{2}}\sqrt{\mathbf{E}_{2}}\dots\sqrt{\mathbf{F}_{n}}\sqrt{\mathbf{E}_{n}}\sqrt{\mathbf{E}_{n}}\sqrt{\mathbf{F}_{n}}\dots\sqrt{\mathbf{E}_{2}}\sqrt{\mathbf{F}_{2}}\sqrt{\mathbf{E}_{1}}\sqrt{\mathbf{F}_{1}}.$ (30)
In some cases, we can take $\mathbf{F}_{1}=\mathbf{F}_{2}=\dots=\mathbf{I}$, and thus $\mathbf{Y}_{n}=\sqrt{\mathbf{E}_{1}}\dots\sqrt{\mathbf{E}_{n}}\sqrt{\mathbf{E}_{n}}\dots\sqrt{\mathbf{E}_{1}}$.
Proof.
Note that
$\mathbb{E}(\mathbf{Y}_{n}|\mathcal{F}_{n-1})=\sqrt{\mathbf{F}_{1}}\sqrt{\mathbf{E}_{1}}\sqrt{\mathbf{F}_{2}}\sqrt{\mathbf{E}_{2}}\dots\sqrt{\mathbf{F}_{n}}\mathbb{E}({\mathbf{E}_{n}}|\mathcal{F}_{n-1})\sqrt{\mathbf{F}_{n}}\dots\sqrt{\mathbf{E}_{2}}\sqrt{\mathbf{F}_{2}}\sqrt{\mathbf{E}_{1}}\sqrt{\mathbf{F}_{1}}$ (31)
$\preceq$ $\sqrt{\mathbf{F}_{1}}\sqrt{\mathbf{E}_{1}}\sqrt{\mathbf{F}_{2}}\sqrt{\mathbf{E}_{2}}\dots\sqrt{\mathbf{F}_{n}}\mathbf{F}_{n}^{-1}\sqrt{\mathbf{F}_{n}}\dots\sqrt{\mathbf{E}_{2}}\sqrt{\mathbf{F}_{2}}\sqrt{\mathbf{E}_{1}}\sqrt{\mathbf{F}_{1}}=\mathbf{Y}_{n-1},$ (32)
concluding the proof. ∎
We remark that by defining $\mathbf{E}_{n}^{\prime}=\sqrt{\mathbf{F}_{n}}\mathbf{E}_{n}\sqrt{\mathbf{F}_{n}}$ and $\mathbf{F}_{n}^{\prime}=\mathbf{I}$, we can obtain another $\mathcal{S}_{d}^{+}$-valued supermartingale $\sqrt{\mathbf{E}_{1}^{\prime}}\dots\sqrt{\mathbf{E}_{n}^{\prime}}\sqrt{\mathbf{E}_{n}^{\prime}}\dots\sqrt{\mathbf{E}_{1}^{\prime}}$. This option, however, involves more computation than (30) which can be computed incrementally by computing the two matrices $\sqrt{\mathbf{F}_{1}}\sqrt{\mathbf{E}_{1}}\sqrt{\mathbf{F}_{2}}\sqrt{\mathbf{E}_{2}}\dots\sqrt{\mathbf{F}_{n}}\sqrt{\mathbf{E}_{n}}$ and $\sqrt{\mathbf{E}_{n}}\sqrt{\mathbf{F}_{n}}\dots\sqrt{\mathbf{E}_{2}}\sqrt{\mathbf{F}_{2}}\sqrt{\mathbf{E}_{1}}\sqrt{\mathbf{F}_{1}}$ on the fly.
With Lemma 4.5, we can recover matrix versions of several recent “test supermartingales” (nonnegative supermartingales with initial value one) in the literature. In the following examples, we will state some conditions on a sequence of $\mathcal{S}_{d}$-valued random matrices $\{\mathbf{X}_{n}\}$ that will lead to a sequence of $\{\mathbf{E}_{n}\}$ and $\{\mathbf{F}_{n}\}$ that satisfy Lemma 4.5, so that a $\mathcal{S}_{d}^{+}$-valued supermartingale $\{\mathbf{Y}_{n}\}$ can be constructed accordingly via (30).
First, we have the MGF-type supermartingales whose well-known scalar case is mentioned by e.g. Howard et al. [2020, Fact 1(a)].
Example 4.6 (Matrix MGF Supermartingale).
Let $\{\mathbf{X}_{n}\}$ be $\mathcal{S}_{d}$-valued random matrices adapted to $\{\mathcal{F}_{n}\}$ with common conditional mean matrix $\mathbb{E}(\mathbf{X}_{n}|\mathcal{F}_{n-1})=\mathbf{M}$ and matrix-valued conditional MGF upper bound $\mathbf{G}(\gamma)\succeq\mathbb{E}(\mathrm{e}^{\gamma(\mathbf{X}_{n}-\mathbf{M})}|\mathcal{F}_{n-1})$. Then, for each $n$, we can take $\mathbf{E}_{n}=\mathrm{e}^{\gamma_{n}(\mathbf{X}_{n}-\mathbf{M})}$, $\mathbf{F}_{n}=\mathbf{G}(\gamma_{n})^{-1}$ where $\{\gamma_{n}\}$ are predictable scalars, in Lemma 4.5.
The MGF upper bounds to use may be found in Table 1. Second, the “betting” style martingales by Waudby-Smith and Ramdas [2023].
Example 4.7 (Betting).
Let $\{\mathbf{X}_{n}\}$ be $\mathcal{S}_{d}^{+}$-valued random matrices adapted to $\{\mathcal{F}_{n}\}$ with conditional means $\mathbb{E}(\mathbf{X}_{n}|\mathcal{F}_{n-1})=\mathbf{M}_{n}$ and upper bounds $\mathbf{B}_{n}$ that are predictable (i.e., $\mathbf{X}_{n}\preceq\mathbf{B}_{n}$ a.s., and $\mathbf{B}_{n}$ is $\mathcal{F}_{n-1}$-measurable). We can take, in Lemma 4.5,
$\mathbf{E}_{n}=\mathbf{I}+\gamma_{n}(\mathbf{X}_{n}-\mathbf{M}_{n}),\quad\mathbf{F}_{n}=\mathbf{I},$ (33)
where
$-\frac{1}{\lambda_{\max}(\mathbf{B}_{n}-\mathbf{M}_{n})}<\gamma_{n}<\frac{1}{\lambda_{\max}(\mathbf{M}_{n})}$ (34)
is predictable.
To see that this is a valid $(\mathbf{E}_{n},\mathbf{F}_{n})$ pair for Lemma 4.5, it is clear that $\mathbb{E}(\mathbf{E}_{n}|\mathcal{F}_{n-1})=\mathbf{I}$ and (34) ensures $\mathbf{E}_{n}\succeq 0$.
Besides these two examples where we extend prior scalar results to matrices, we also note that Howard et al. [2020, Section 6.5] prove a total of eight inequalities for matrix martingale differences $\{\mathbf{Z}_{n}\}$ in the style of
$\mathbb{E}(\exp(\mathbf{Z}_{n}-\mathbf{C}_{n})|\mathcal{F}_{n-1})\preceq\exp(\mathbf{C}_{n}^{\prime}),$ (35)
for some respectively $\mathcal{F}_{n}$-, $\mathcal{F}_{n-1}$-measurable matrices $\mathbf{C}_{n},\mathbf{C}_{n}^{\prime}$ under eight different conditions. These directly lead to $\mathcal{S}_{d}^{+}$-valued supermartingales via Lemma 4.5. However, Howard et al. [2020] do not consider combining these conditional expectation bounds into matrix-valued supermartingales via (30). For reasons we will discuss later in Section 5; they pursue scalar supermartingales instead. Of their eight inequalities, we will introduce two interesting cases now, omitting the other six due to their similar construction.
First, the self-normalized supermartingale that only assumes the second moment exists [Howard et al., 2020, Part (f) of Section 6.5], which we (re)prove in Section B.3.
Example 4.8 (Self-Normalized Matrix Supermartingale).
Let $\{\mathbf{X}_{n}\}$ be $\mathcal{S}_{d}$-valued random matrices adapted to $\{\mathcal{F}_{n}\}$ with conditional means $\mathbb{E}(\mathbf{X}_{n}|\mathcal{F}_{n-1})=\mathbf{M}_{n}$ and conditional variance upper bounds $\mathbb{E}((\mathbf{X}_{n}-\mathbf{M}_{n})^{2}|\mathcal{F}_{n-1})\preceq\mathbf{V}_{n}$ that are predictable. We can take, in Lemma 4.5,
$\mathbf{E}_{n}=\exp(\gamma_{n}(\mathbf{X}_{n}-\mathbf{M}_{n})-\gamma_{n}^{2}(\mathbf{X}_{n}-\mathbf{M}_{n})^{2}/6),\quad\mathbf{F}_{n}=\exp\left(-\frac{\gamma_{n}^{2}}{3}\mathbf{V}_{n}\right)$ (36)
where $\{\gamma_{n}\}$ are predictable positive scalars.
Second, a bound for symmetric random matrices that have a symmetric (around $0\in\mathcal{S}_{d}$), potentially heavy-tailed distribution [Howard et al., 2020, Part (d) of Section 6.5] which we prove as well in Section B.3.
Example 4.9.
Let $\{\mathbf{X}_{n}\}$ be $\mathcal{S}_{d}$-valued random matrices adapted to $\{\mathcal{F}_{n}\}$ with conditional means $\mathbb{E}(\mathbf{X}_{n}|\mathcal{F}_{n-1})=\mathbf{M}_{n}$ and satisfy $\mathbf{X}_{n}-\mathbf{M}_{n}|\mathcal{F}_{n-1}\stackrel{{\scriptstyle d}}{{=}}\mathbf{M}_{n}-\mathbf{X}_{n}|\mathcal{F}_{n-1}$. Then, we can take, in Lemma 4.5,
$\mathbf{E}_{n}=\exp\left(\gamma_{n}(\mathbf{X}_{n}-\mathbf{M}_{n})-\frac{\gamma_{n}^{2}(\mathbf{X}_{n}-\mathbf{M}_{n})^{2}}{2}\right),\quad\mathbf{F}_{n}=\mathbf{I},$ (37)
where $\{\gamma_{n}\}$ are predictable positive scalars.
Finally, let us present a novel matrix extension of the empirical Bernstein bound, in the style of Howard et al. [2021, Theorem 4], following whom we define the function
$\psi_{E}(x)=-\log(1-x)-x=\sum_{k=2}^{\infty}\frac{x^{k}}{k}$ (38)
for $x\in[0,1)$. The notation comes from the fact that $\psi_{E}$ is the log-MGF of a centered standard exponential random variable. We distinguish the following result for its novelty by labeling it a “proposition” as opposed to an “example”.
Proposition 4.10 (Empirical Bernstein Matrix Supermartingale).
Let $\{\mathbf{X}_{n}\}$ be $\mathcal{S}_{d}$-valued random matrices adapted to $\{\mathcal{F}_{n}\}$ with conditional means $\mathbb{E}(\mathbf{X}_{n}|\mathcal{F}_{n-1})=\mathbf{M}_{n}$. Further, suppose there is a predictable and integrable sequence of $\mathcal{S}_{d}$-valued random matrices $\{\widehat{\mathbf{X}}_{n}\}$ such that $\lambda_{\min}(\mathbf{X}_{n}-\widehat{\mathbf{X}}_{n})\geqslant-1$. Then, we can take, in Lemma 4.5,
$\mathbf{E}_{n}=\exp(\gamma_{n}(\mathbf{X}_{n}-\widehat{\mathbf{X}}_{n})-\psi_{E}(\gamma_{n})(\mathbf{X}_{n}-\widehat{\mathbf{X}}_{n})^{2}),\quad\mathbf{F}_{n}=\exp(-\gamma_{n}(\mathbf{M}_{n}-\widehat{\mathbf{X}}_{n})),$ (39)
where $\{\gamma_{n}\}$ are predictable $(0,1)$-valued scalars.
We delay the proof that $\mathbb{E}(\mathbf{E}_{n}|\mathcal{F}_{n-1})\preceq\mathbf{F}_{n}^{-1}$ to Section B.4. The reason why it is called “empirical” Bernstein is the “plug-in empirical variance” component term $(\mathbf{X}_{n}-\widehat{\mathbf{X}}_{n})^{2}$, as $\widehat{\mathbf{X}}_{n}$ can usually be taken as an empirical average (possibly weighted) from $\mathbf{X}_{1}$ to $\mathbf{X}_{n-1}$, in which case the condition $\lambda_{\min}(\mathbf{X}_{n}-\widehat{\mathbf{X}}_{n})\geqslant-1$ is satisfied when $\{\mathbf{X}_{n}\}$ all take values in $\mathcal{S}_{d}^{[0,1]}$. We shall come to this empirical Bernstein result again with some scalarization tricks in Section 5.3, where we obtain a strikingly sharp spectral concentration inequality in closed form, further justify its name, discuss the context of empirical Bernstein bounds, and compare it with other existing results.
These supermartingales lead to implicit concentration bounds when applied Theorem 4.4, and are thus unfavorable when one needs a confidence interval or confidence sequence. However, these supermartingales can be used to test whether $\mathbb{E}\mathbf{X}_{1}=\mathbf{M}$ or not and are thus “test supermartingales” only.
4.3 Matrix e-Processes
Scalar e-processes are an important class of processes generalizing scalar nonnegative supermartingales [Ramdas et al., 2022]. We recall the definition of these processes below, noting that we will always consider a sequence of observations $X_{1},X_{2},\dots$ (while allowing their distribution to vary) and denote their canonical filtration by $\{\mathcal{F}_{n}\}$. We use notations $\mathbb{P}_{P}$, $\mathbb{E}_{P}$ and the general terminology “for $P$” (e.g. we may say a stochastic process adapted to $\{\mathcal{F}_{n}\}$ is a martingale for $P$) to refer to probabilistic statements when the observations $X_{1},X_{2},\dots$ are distributed jointly according to $P$.
Let $\mathcal{P}$ be a set of distributions over $\mathbb{R}\times\mathbb{R}\times\dots$. A nonnegative stochastic process $\{Y_{n}\}$ adapted to $\{\mathcal{F}_{n}\}$ is called an e-process for $\mathcal{P}$ if, for any $P\in\mathcal{P}$, there exists a supermartingale $\{M_{n}^{P}\}$ adapted to $\{\mathcal{F}_{n}\}$ with $M_{0}^{P}\leqslant 1$ that upper bounds the process $\{Y_{n}\}$. Equivalently, $\{Y_{n}\}$ is an e-process if $\mathbb{E}_{P}[Y_{\tau}]\leqslant 1$ for any stopping time $\tau$ on $\{\mathcal{F}_{n}\}$ [Ramdas et al., 2020, Lemma 6]. Clearly, if $\{Y_{n}\}$ is a nonnegative supermartingale for $\mathcal{P}$ (meaning for every $P\in\mathcal{P}$), the optional stopping theorem implies that it is also an e-process for $\mathcal{P}$, but e-processes are a much more general class [Ramdas et al., 2022, Ruf et al., 2023].
A scalar e-process satisfies the following form of (VI) or (UVI):
for $\text{for any }P\in\mathcal{P},\mathbb{P}_{P}(Y_{\tau}\geqslant Ua)\leqslant a^{-1}$
for any stopping time $\tau$ and independent super-uniform random variable $U$ (and thus we also have that $\mathbb{P}_{P}(Y_{\tau}\geqslant a)\leqslant a^{-1}$ by setting $U=1$). This property enables scalar e-processes to be used for testing a hypothesis that the data is drawn from some $P\in\mathcal{P}$: we may reject the null as soon as the e-process exceeds $1/\alpha$ (for some prespecified constant $\alpha\in(0,1)$) and be guaranteed that the type-I error is at most $\alpha$. E-processes are particularly useful for composite sequential testing problems for which nonnegative supermartingales are powerless [Ramdas et al., 2022, Ruf et al., 2023].
We can generalize the scalar e-processes to our matrix setting as well, with the sequence of observations that generates the filtration $\{\mathcal{F}_{n}\}$ now denoted by $\mathbf{X}_{1},\mathbf{X}_{2},\dots$.
Definition 4.11 (Matrix e-Process).
Let $\mathfrak{P}$ be a set of distributions over $\mathcal{S}_{d}\times\mathcal{S}_{d}\times\dots$. An $\mathcal{S}_{d}^{+}$-valued stochastic process $\{\mathbf{Y}_{n}\}$ adapted to $\{\mathcal{F}_{n}\}$ is called a matrix e-process for $\mathfrak{P}$ if, for any $\mathfrak{p}\in\mathfrak{P}$, there exists an $\mathcal{S}_{d}^{+}$-valued supermartingale $\{\mathbf{M}_{n}^{\mathfrak{p}}\}$ adapted to $\{\mathcal{F}_{n}\}$ with $\mathbf{M}_{0}^{\mathfrak{p}}\preceq\mathbf{I}$ that upper bounds the process $\{\mathbf{Y}_{n}\}$, i.e., $\mathbf{M}_{n}^{\mathfrak{p}}\succeq\mathbf{Y}_{n}$ for all $n$ almost surely for $\mathfrak{p}$.
Similar to the scalar case, matrix e-processes satisfy their (UMVI).
Lemma 4.12 (Uniformly Randomized Matrix Ville’s Inequality for Matrix e-Processes).
Let $\{\mathbf{Y}_{n}\}$ be a matrix e-process for $\mathfrak{P}$. Then, for any $\mathfrak{p}\in\mathfrak{P}$, for any stopping time $\tau$ on $\{\mathcal{F}_{n}\}$, any trace super-uniform random matrix $\mathbf{U}$ taking values in $\mathcal{S}_{d}^{+}$ independent from $\mathbf{X}_{1},\mathbf{X}_{2},\dots$ and any $\mathbf{A}\in\mathcal{S}_{d}^{++}$,
$\mathbb{P}_{\mathfrak{p}}(\mathbf{Y}_{\tau}\npreceq\mathbf{A}^{1/2}\mathbf{U}\mathbf{A}^{1/2})\leqslant\mathrm{tr}(\mathbf{A}^{-1}).$ (40)
Proof.
Consider the upper bounding matrix supermartingale $\{\mathbf{M}_{n}^{\mathfrak{p}}\}$. By (UMVI), $\mathbb{P}(\mathbf{M}_{\tau}^{\mathfrak{p}}\npreceq\mathbf{A}^{1/2}\mathbf{U}\mathbf{A}^{1/2})\leqslant\mathrm{tr}(\mathbf{A}^{-1})$. Since $\mathbf{M}_{n}^{\mathfrak{p}}\succeq\mathbf{Y}_{n}$ almost surely for any $n$, $\mathbf{M}_{\tau}^{\mathfrak{p}}\succeq\mathbf{Y}_{\tau}$ almost surely as well. Therefore, $\mathbb{P}(\mathbf{Y}_{\tau}\npreceq\mathbf{A}^{1/2}\mathbf{U}\mathbf{A}^{1/2})\leqslant\mathbb{P}(\mathbf{M}_{\tau}^{\mathfrak{p}}\npreceq\mathbf{A}^{1/2}\mathbf{U}\mathbf{A}^{1/2})\leqslant\mathrm{tr}(\mathbf{A}^{-1}).$
Simple examples of matrix e-processes following the previous subsection include lower bounds of $\mathcal{S}_{d}^{+}$-valued supermartingales. First, we define the following “matrix minimum”: for $\mathbf{A},\mathbf{B}\in\mathcal{S}_{d}$, we define $\mathbf{A}\curlyvee\mathbf{B}$ to be $\mathbf{A}$ if $\mathbf{A}\preceq\mathbf{B}$, and $\mathbf{A}-\lambda_{\max}(\mathbf{A}-\mathbf{B})\cdot\mathbf{I}$ if otherwise. Note that $\mathbf{A}\curlyvee\mathbf{B}\preceq\mathbf{A}$ and $\mathbf{A}\curlyvee\mathbf{B}\preceq\mathbf{B}$, and we remark that in general there is no largest symmetric matrix lower bound of $\mathbf{A}$ and $\mathbf{B}$ in the $\preceq$ order [Kadison, 1951, Theorem 6].
Let $\mathbf{M},\mathbf{M}^{\prime}\preceq\mathbf{B}$ be three matrices in $\mathcal{S}_{d}^{++}$, and we define $\mathcal{B}(\mathbf{B},\mathbf{M})$ to be the set of distributions on $\mathcal{S}_{d}^{+}$ upper bounded by $\mathbf{B}$ with mean $\mathbf{M}$. Further, we define the distribution families
$\mathfrak{P}=\{P_{1}\otimes P_{2}\otimes\dots:P_{n}\in\mathcal{B}(\mathbf{B},\mathbf{M})\text{ for all }n\}$ (41)
and
$\mathfrak{P}^{\prime}=\{P_{1}\otimes P_{2}\otimes\dots:P_{n}\in\mathcal{B}(\mathbf{B},\mathbf{M}^{\prime})\text{ for all }n\}.$ (42)
According to Example 4.7, we can construct a $\mathcal{S}_{d}^{+}$-valued betting process $\{\mathbf{Y}_{n}\}$ which is a supermartingale under any of $\mathfrak{p}\in\mathfrak{P}$, and another $\{\mathbf{Y}_{n}^{\prime}\}$ which is a supermartingale under any of $\mathfrak{p}\in\mathfrak{P}^{\prime}$. Then the minimum process $\{\mathbf{Y}_{n}\curlyvee\mathbf{Y}_{n}^{\prime}\}$ is a matrix $\mathfrak{P}\cup\mathfrak{P}^{\prime}$-e-process, useful for testing sequentially the null that the true mean is always either $\mathbf{M}$ or $\mathbf{M}^{\prime}$.
4.4 Matrix-Valued Submartingales and Backward Supermartingales
Next, we note that all the above are still valid if we replace “supermartingale”s with “submartingale”s. To wit, we say that $\{\mathbf{Y}_{n}\}$ is a submartingale if $\mathbb{E}(\mathbf{Y}_{n}|\mathcal{F}_{n-1})\succeq\mathbf{Y}_{n-1}$. It is clear from the proof of Lemma 4.3 that,
Lemma 4.13.
Let $\{\mathbf{Y}_{n}\}$ be an $\mathcal{S}_{d}$-valued submartingale, and $\tau$ be a stopping time on $\{\mathcal{F}_{n}\}$. Then, $\{\mathbf{Y}_{n\wedge\tau}\}$ is also a submartingale on $\{\mathcal{F}_{n}\}$. Consequently, if $\tau\leqslant N$ almost surely for some $N\in\mathbb{N}$, then $\mathbb{E}\mathbf{Y}_{\tau}\preceq\mathbb{E}\mathbf{Y}_{N}$.
Doob’s submartingale inequality, which is “dual” to Ville’s supermartingale inequality, is also matrix generalizable.
Theorem 4.14 (Matrix Doob’s inequality).
Let $\{\mathbf{Y}_{n}\}$ be an $\mathcal{S}_{d}^{+}$-valued submartingale. For every $N\in\mathbb{N}$ and every $\mathbf{A}\in\mathcal{S}_{d}^{++}$,
$\mathbb{P}\left(\exists n\leqslant N,\mathbf{Y}_{n}\npreceq\mathbf{A}\right)\leqslant\mathrm{tr}(\mathbb{E}(\mathbf{Y}_{N}\mathbbmss{1}_{\{\exists n\leqslant N,\mathbf{Y}_{n}\npreceq\mathbf{A}\}})\mathbf{A}^{-1})\leqslant\mathrm{tr}((\mathbb{E}\mathbf{Y}_{N})\mathbf{A}^{-1}).$ (43)
Proof.
Let $E$ be the event $\{\exists n\leqslant N,\mathbf{Y}_{n}\npreceq\mathbf{A}\}$. Define the stopping time
$\tau=\inf\{n:\mathbf{Y}_{n}\npreceq\mathbf{A}\}\wedge N,$ (44)
so it satisfies $\tau\leqslant N$ and
$\{\mathbf{Y}_{\tau}\mathbbmss{1}_{E}\npreceq\mathbf{A}\}=E,\quad\mathbf{Y}_{\tau}\mathbbmss{1}_{\overline{E}}=\mathbf{Y}_{N}\mathbbmss{1}_{\overline{E}}.$ (45)
Thus by Lemmas 4.13 and 3.3,
$\mathrm{tr}(\mathbf{A}^{-1/2}(\mathbb{E}\mathbf{Y}_{N})\mathbf{A}^{-1/2})\geqslant\mathrm{tr}(\mathbf{A}^{-1/2}(\mathbb{E}\mathbf{Y}_{\tau})\mathbf{A}^{-1/2})$ (46)
$=$ $\mathrm{tr}(\mathbf{A}^{-1/2}\mathbb{E}(\mathbf{Y}_{\tau}\mathbbmss{1}_{E})\mathbf{A}^{-1/2})+\mathrm{tr}(\mathbf{A}^{-1/2}\mathbb{E}(\mathbf{Y}_{\tau}\mathbbmss{1}_{\overline{E}})\mathbf{A}^{-1/2})$ (47)
$\geqslant$ $\mathbb{P}(\mathbf{Y}_{\tau}\mathbbmss{1}_{E}\npreceq\mathbf{A})+\mathrm{tr}(\mathbf{A}^{-1/2}\mathbb{E}(\mathbf{Y}_{N}\mathbbmss{1}_{\overline{E}})\mathbf{A}^{-1/2}),$ (48)
implying that $\mathbb{P}(E)\leqslant\mathrm{tr}(\mathbf{A}^{-1/2}\mathbb{E}(\mathbf{Y}_{N}\mathbbmss{1}_{E})\mathbf{A}^{-1/2})$. ∎
Next, we illustrate how Theorem 4.14 remains valid even without assuming that the process $\{\mathbf{Y}_{n}\}$ starts at $n=0$ — or starts at all; it is applicable even if $\{\mathbf{Y}_{n}\}$ is indexed on $n\in\{\dots,-2,-1,0,1,\dots,N\}$. We formalize this by considering a backward filtration $\{\mathcal{G}_{n}\}$, i.e., $\mathcal{G}_{0}\supseteq\mathcal{G}_{1}\supseteq\dots$, and an adapted backward submartingale $\{\mathbf{Y}_{n}\}$, i.e., $\mathbf{Y}_{n}$ is $\mathcal{G}_{n}$-measurable and $\mathbb{E}[\mathbf{Y}_{n}|\mathcal{G}_{n+1}]\succeq\mathbf{Y}_{n+1}$, while keeping the time indices $n\geqslant 0$. This, in the scalar regime, sets the stage for the backward Ville’s inequality which has been shown equivalent to the (forward) Doob’s inequality (hence the duality between Ville’s and Doob’s inequality; see the second proof of Theorem 2 in Manole and Ramdas [2023]), for which we have the following matrix version.
Theorem 4.15 (Matrix Backward Ville’s Inequality).
Let $\{\mathbf{Y}_{n}\}$ be an $\mathcal{S}_{d}^{+}$-valued backward submartingale adapted to the backward filtration $\{\mathcal{G}_{n}\}$. For any $\mathbf{A}\in\mathcal{S}_{d}^{++}$,
$\mathbb{P}(\exists n,\;\mathbf{Y}_{n}\npreceq\mathbf{A})\leqslant\mathrm{tr}((\mathbb{E}\mathbf{Y}_{0})\mathbf{A}^{-1}).$ (49)
Proof.
Take an arbitrary $N\in\mathbb{N}$ and consider, indexed by $0\leqslant n\leqslant N$, the process $\mathbf{P}_{n}=\mathbf{Y}_{N-n}$ on the (forward) filtration $\mathcal{F}_{n}=\mathcal{G}_{N-n}$. It is clear that $\{\mathbf{P}_{n}\}$ is a (forward) submartingale on $\{\mathcal{F}_{n}\}$, to which we can apply Theorem 4.14,
$\mathbb{P}\left(\exists n\leqslant N,\mathbf{P}_{n}\npreceq\mathbf{A}\right)\leqslant\mathrm{tr}((\mathbb{E}\mathbf{P}_{N})\mathbf{A}^{-1}).$ (50)
Substituting $M_{n}$ back to the inequality above, we have,
$\mathbb{P}\left(\exists n\leqslant N,\mathbf{Y}_{n}\npreceq\mathbf{A}\right)\leqslant\mathrm{tr}((\mathbb{E}\mathbf{Y}_{0})\mathbf{A}^{-1}).$ (51)
The desired inequality is immediately proved by appealing to the fact that
$\{\exists n,\;\mathbf{Y}_{n}\npreceq\mathbf{A}\}=\liminf_{N\to\infty}\left\{\exists n\leqslant N,\;\mathbf{Y}_{n}\npreceq\mathbf{A}\right\}=\bigcup_{k=1}^{\infty}\bigcap_{N=k}^{\infty}\left\{\exists n\leqslant N,\;\mathbf{Y}_{n}\npreceq\mathbf{A}\right\}.$ (52)
4.5 Exchangeable Matrices
We now have the tools to prove a matrix version of the exchangeable Chebyshev inequality (XCI) we mentioned earlier. Recall from Fact 2 that when $p\in[1,2]$, the function $\mathbf{X}\mapsto\mathbf{X}^{p}$ is operator convex within the domain $\mathcal{S}_{d}^{+}$ (for $21\leqslant p<$ 2) or $\mathcal{S}_{d}$ (for $p=2$) respectively,
$\frac{1}{n}(\mathbf{X}_{1}^{p}+\dots+\mathbf{X}_{n}^{p})\succeq\left(\frac{\mathbf{X}_{1}+\dots+\mathbf{X}_{n}}{n}\right)^{p}.$ (53)
This gives rise to the following backward submartingale for exchangeable matrices.
Lemma 4.16.
Let $p\in[1,2]$, and $\{\mathbf{X}_{n}\}$ be an exchangeable sequence of $\mathcal{S}_{d}$-valued, $L^{p}$ random matrices. Suppose $\mathbf{C}\in\mathcal{S}_{d}$ is deterministic, and either of the following holds:
  • $p=2$; or
  • $\mathbf{X}_{n}\succeq\mathbf{C}$ almost surely for all $n$.
Then, the process
$\mathbf{R}_{n}=(\overline{\mathbf{X}}_{n}-\mathbf{C})^{p}$ (54)
is a backward submartingale adapted to the exchangeable backward filtration
$\mathcal{E}_{n}=\sigma(f(\mathbf{X}_{1},\mathbf{X}_{2},\dots):f\text{ symmetric w.r.t.\ first $n$ arguments}).$ (55)
Proof.
Recalling that the conditional expectation w.r.t. the exchangeable $\sigma$-algebra $\mathcal{E}_{n+1}$ equals the average over all $(n+1)!$ permutations of the indices $11,2,\dots,n+$ 1 (see e.g. Klenke [2013, Theorem 12.10]), we have
$\mathbb{E}[\mathbf{R}_{n}|\mathcal{E}_{n+1}]=\frac{1}{(n+1)!}\sum_{\sigma\in\mathrm{S}_{n+1}}\left(\frac{\mathbf{X}_{\sigma(1)}+\dots+\mathbf{X}_{\sigma(n)}}{n}-\mathbf{C}\right)^{p}$ (56)
(${{\succeq}}$ $\left(\frac{\sum_{\sigma\in\mathrm{S}_{n+1}}\left(\frac{\mathbf{X}_{\sigma(1)}+\dots+\mathbf{X}_{\sigma(n)}}{n}-\mathbf{C}\right)}{(n+1)!}\right)^{p}=\mathbf{R}_{n+1},$ (57)
concluding the proof that $\{\mathbf{R}_{n}\}$ is a backward submartingale. ∎
Applying the matrix backward Ville’s inequality (Theorem 4.15) to this process $\{\mathbf{R}_{n}\}$, we immediately have the following two matrix extensions of (XCI).
Theorem 4.17 (Exchangeable Matrix Chebyshev Inequality).
Let $\{\mathbf{X}_{n}\}$ be an exchangeable sequence of $\mathcal{S}_{d}$-valued random matrices with mean $\mathbf{M}$ and variance upper bound $\mathbb{E}(\mathbf{X}_{1}-\mathbf{M})^{2}\preceq\mathbf{V}$. Then, for any $\mathbf{A}\in\mathcal{S}_{d}^{++}$,
$\mathbb{P}(\exists n,\;\operatorname{abs}(\overline{\mathbf{X}}_{n}-\mathbf{M})\npreceq\mathbf{A})\leqslant\mathrm{tr}(\mathbf{V}\mathbf{A}^{-2}).$ (XMCI)
Further, if $\mathbb{E}\{\mathbf{X}_{i}-\mathbf{M},\mathbf{X}_{j}-\mathbf{M}\}=0$ for any $i\neq j$, we have for any fixed $N$,
$\mathbb{P}(\exists n\geqslant N,\;\operatorname{abs}(\overline{\mathbf{X}}_{n}-\mathbf{M})\npreceq\mathbf{A})\leqslant N^{-1}\mathrm{tr}(\mathbf{V}\mathbf{A}^{-2}).$ (58)
Proof.
Applying Theorem 4.15 to $\mathbf{R}_{n}=(\overline{\mathbf{X}_{n}}-\mathbf{M})^{2}$, which is a backward submartingale according to Lemma 4.16,
$\mathbb{P}(\exists n,\;(\overline{\mathbf{X}}_{n}-\mathbf{M})^{2}\npreceq\mathbf{A}^{2})\leqslant\mathrm{tr}((\mathbb{E}\mathbf{R}_{1})\mathbf{A}^{-2})\leqslant\mathrm{tr}(\mathbf{V}\mathbf{A}^{-2}).$ (59)
Now use the operator monotonicity of $\mathbf{X}\mapsto\mathbf{X}^{1/2}$,
$\mathbb{P}(\exists n,\;\operatorname{abs}(\overline{\mathbf{X}}_{n}-\mathbf{M})\npreceq\mathbf{A})\leqslant\mathbb{P}(\exists n,\;(\overline{\mathbf{X}}_{n}-\mathbf{M})^{2}\npreceq\mathbf{A}^{2}),$ (60)
concluding the proof of (XMCI). To show (58), simply apply Theorem 4.15 to the submartingale $\mathbf{R}_{n}=(\overline{\mathbf{X}_{n}}-\mathbf{M})^{2}$ with $n\geqslant N$, obtaining
$\mathbb{P}(\exists n\geqslant N,\;(\overline{\mathbf{X}}_{n}-\mathbf{M})^{2}\npreceq\mathbf{A}^{2})\leqslant\mathrm{tr}((\mathbb{E}\mathbf{R}_{n})\mathbf{A}^{-2})\leqslant N^{-1}\mathrm{tr}(\mathbf{V}\mathbf{A}^{-2})$ (61)
due to Theorem 3.5. ∎
Theorem 4.18 (Exchangeable Matrix $p$-Chebyshev Inequality).
Let $\{\mathbf{X}_{n}\}$ be an exchangeable sequence of $\mathcal{S}_{d}^{+}$-valued random matrices with raw $p$th moment $\mathbf{V}_{p}=\mathbb{E}\mathbf{X}_{1}^{p}$, where $p\in[1,2]$. Then, for any $\mathbf{A}\in\mathcal{S}_{d}^{++}$,
$\mathbb{P}(\exists n,\;\overline{\mathbf{X}}_{n}\npreceq\mathbf{A})\leqslant\mathrm{tr}(\mathbf{V}_{p}\mathbf{A}^{-p}).$ (62)
Proof.
Applying Theorem 4.15 to $\{\mathbf{R}_{n}\}$, with $\mathbf{C}$ being 00, in Lemma 4.16,
$\mathbb{P}(\exists n,\;(\overline{\mathbf{X}}_{n})^{p}\npreceq\mathbf{A}^{p})\leqslant\mathrm{tr}(\mathbf{V}_{p}\mathbf{A}^{-p}).$ (63)
Now use the operator monotonicity of $\mathbf{X}\mapsto\mathbf{X}^{1/p}$,
$\mathbb{P}(\exists n,\;\overline{\mathbf{X}}_{n}\npreceq\mathbf{A})\leqslant\mathbb{P}(\exists n,\;(\overline{\mathbf{X}}_{n})^{p}\npreceq\mathbf{A}^{p}),$ (64)
concluding the proof. ∎
Indeed, we can weaken the second-moment assumption to a $p$th moment, leading to, however, a concentration not around the mean but a Markov-type upper bound. This is largely because the scalar map $x\mapsto|x|^{p}$ is convex but not its matrix extension, and so the raw $p$th moment is used above. Nonetheless, due to Fact 4, the following trace inequality holds.
$\mathrm{tr}\left(\frac{1}{n}((\operatorname{abs}\mathbf{X}_{1})^{p}+\dots+(\operatorname{abs}\mathbf{X}_{n})^{p})\right)\geqslant\mathrm{tr}\left(\left(\operatorname{abs}\frac{\mathbf{X}_{1}+\dots+\mathbf{X}_{n}}{n}\right)^{p}\right).$ (65)
Therefore,
$r_{n}=\mathrm{tr}\left(\operatorname{abs}(\overline{\mathbf{X}}_{n}-\mathbf{C})^{p}\right)$ (66)
is a real-valued backward submartingale. With the scalar backward Ville’s inequality, this leads to the following anytime-valid extension of Lemma 3.7.
Theorem 4.19 (Exchangeable Matrix $p$-Chebyshev Trace Inequality).
Let $\{\mathbf{X}_{n}\}$ be an exchangeable sequence of $\mathcal{S}_{d}$-valued random matrices with mean $\mathbf{M}$ and central $p$th moment $\mathbf{V}_{p}=\mathbb{E}\operatorname{abs}(\mathbf{X}_{1}-\mathbf{M})^{p}$, where $p\geqslant 1$. Then, for any $a>0$,
$\mathbb{P}(\exists n,\;\|\overline{\mathbf{X}}_{n}-\mathbf{M}\|_{\mathrm{op}}\geqslant a)\leqslant\mathbb{P}\left(\exists n,\;\mathrm{tr}\left(\operatorname{abs}(\overline{\mathbf{X}}_{n}-\mathbf{M})^{p}\right)\geqslant a^{p}\right)\leqslant a^{-p}\mathrm{tr}\mathbf{V}_{p}.$ (67)
The bound above differs from Theorem 4.18 in that the concentration is now around the mean $\mathbf{M}$.
5 Randomized Matrix Bounds via Scalar e-Processes
5.1 A Conversion Lemma from Matrix to Scalar Supermartingales
We have studied the concentration of random matrices via matrix e-statistics, i.e. random matrices with bounded expectations, matrix supermartingales, matrix e-values, and matrix backward submartingales, since these quantities yield concentration inequalities via our Markov and Ville’s inequalities for matrices, oftentimes in randomized forms. Some further randomized matrix concentration inequalities can be obtained by revisiting some bounds obtained via scalar e-quantities that already exist in the literature.
As we mentioned earlier in Section 4.2, Howard et al. [2020, Section 6.5] prove, under different conditions, numerous inequalities for symmetric matrix martingale difference $\{\mathbf{Z}_{n}\}$ in the style of
$\mathbb{E}(\exp(\mathbf{Z}_{n}-\mathbf{C}_{n})|\mathcal{F}_{n-1})\preceq\exp(\mathbf{C}_{n}^{\prime}),$ (68)
for some respectively $\mathcal{F}_{n}$-, $\mathcal{F}_{n-1}$-measurable symmetric matrices $\mathbf{C}_{n},\mathbf{C}_{n}^{\prime}$.
At this step, our previous methods and that of Howard et al. [2020] had diverged. We have been able to directly construct matrix-valued supermartingales in Section 4.2, leading potentially to computable sequential tests by virtue of our new (UMVI). Howard et al. [2020], on the other hand, construct scalar e-processes from the condition (68). Let us rephrase their “master lemma” behind this construction.
Lemma 5.1 (Lemma 4 in Howard et al. [2020], rephrased and generalized).
Let $\{\mathbf{Z}_{n}\}$ be an $\mathcal{S}_{d}$-valued martingale difference sequence adapted to $\{\mathcal{F}_{n}\}$. Let $\{\mathbf{C}_{n}\}$ be an $\mathcal{S}_{d}$-valued adapted process, $\{\mathbf{C}_{n}^{\prime}\}$ be an $\mathcal{S}_{d}$-valued predictable process, w.r.t. the same filtration. If (68) holds for all $n$, then, the process
$L_{n}=\mathrm{tr}\exp\left(\sum_{i=1}^{n}\mathbf{Z}_{i}-\sum_{i=1}^{n}(\mathbf{C}_{i}+\mathbf{C}_{i}^{\prime})\right)$ (69)
is a (scalar) supermartingale. Further,
$L_{n}\geqslant\exp\left(\lambda_{\max}\left(\sum_{i=1}^{n}\mathbf{Z}_{i}\right)-\lambda_{\max}\left(\sum_{i=1}^{n}(\mathbf{C}_{i}+\mathbf{C}_{i}^{\prime})\right)\right).$ (70)
The proof of this lemma can be found in Section B.5. Note that the upper bounded expression (70) is an instance of e-processes, introduced earlier in Section 4.3. We shall see soon that both (69) and (70) can be useful.
5.2 Examples: Self-Normalization, Bernstein, and Hoeffding
Recall from Example 4.8 the self-normalized bound, that if $\{\mathbf{X}_{n}\}$ are symmetric matrices adapted to $\{\mathcal{F}_{n}\}$ with conditional means $\{\mathbf{M}_{n}\}$ and predictable conditional variance upper bounds $\{\mathbf{V}_{n}\}$,
$\mathbb{E}\left(\exp(\gamma_{n}(\mathbf{X}_{n}-\mathbf{M}_{n})-\gamma_{n}^{2}(\mathbf{X}_{n}-\mathbf{M}_{n})^{2}/6)\mid\mathcal{F}_{n-1}\right)\preceq\exp\left(\frac{\gamma_{n}^{2}}{3}\mathbf{V}_{n}\right),$ (71)
an instance of (68). By Lemma 5.1, this implies the following statement whose weaker version is stated as Lemma 3(f) by Howard et al. [2020].
Example 5.2 (Matrix Self-Normalized Scalar Supermartingale).
Under the same assumptions as Example 4.8, the process
$L_{n}^{\textsf{SN}}=\mathrm{tr}\exp\left(\sum_{i=1}^{n}\gamma_{i}(\mathbf{X}_{i}-\mathbf{M}_{i})-\sum_{i=1}^{n}\frac{\gamma_{i}^{2}}{6}((\mathbf{X}_{i}-\mathbf{M}_{i})^{2}+2\mathbf{V}_{i})\right)$ (72)
is a supermartingale.
The proof is simply taking $\mathbf{Z}_{n}=\gamma_{n}(\mathbf{X}_{n}-\mathbf{M}_{n})$, $\mathbf{C}_{n}=\frac{\gamma_{n}^{2}}{6}(\mathbf{X}_{n}-\mathbf{M}_{n})^{2}$, and $\mathbf{C}_{n}^{\prime}=\frac{\gamma_{n}^{2}}{3}\mathbf{V}_{n}$ in Lemma 5.1. Some remarks on this result are in order. First, the expression of the scalar supermartingale $\{L_{n}^{\textsf{SN}}\}$ explains the name “self-normalized”, as an empirical variance term exists within the bound. Second, unlike the scalar case where a closed-form confidence sequence can be solved from the supermartingale combining (VI) if the means are a constant [Wang and Ramdas, 2023a, Section 5], in our matrix case, we only have an implicit bound when applying (UVI).
Corollary 5.3 (Uniformly Randomized Stopped Matrix Self-Normalized Bound).
Under the same assumptions as Example 4.8, for any stopping time $\tau$, $\alpha\in(0,1)$, and a super-uniform $U$ independent from $\mathcal{F}_{\infty}$,
$\mathbb{P}\left(\mathrm{tr}\exp\left(\sum_{i=1}^{\tau}\gamma_{i}(\mathbf{X}_{i}-\mathbf{M}_{i})-\sum_{i=1}^{\tau}\frac{\gamma_{i}^{2}}{6}((\mathbf{X}_{i}-\mathbf{M}_{i})^{2}+2\mathbf{V}_{i})\right)\geqslant dU/\alpha\right)\leqslant\alpha.$ (73)
The bound above, as those matrix supermartingales in Section 4.2, does not bear closed-form confidence sets but still implies a computable randomized sequential test for the null, e.g., that the means $\mathbb{E}(\mathbf{X}_{n}|\mathcal{F}_{n-1})$ all equal $\mathbf{M}$ with marginal type I error rate within $\alpha$.
However, a closed-form concentration can be computed when further assumption is introduced on top of the finite variance in Example 4.8. This is the Hoeffding-type e-process by Howard et al. [2020, Lemma 3], which we present as a Bernstein-type result akin to Mackey et al. [2014, Corollary 4.2] for generality.
Example 5.4 (Matrix Bernstein Scalar e-Process).
Under the same assumptions as Example 4.8, suppose further that $(\mathbf{X}_{n}-\mathbf{M}_{n})^{2}\preceq\mathbf{B}_{n}$ almost surely for some predictable sequence $\{\mathbf{B}_{n}\}$. Then, the self-normalized supermartingale $\{L_{n}^{\textsf{SN}}\}$ upper bounds the following process:
$L_{n}^{\textsf{SN}}\geqslant\exp\left\{\lambda_{\max}\left(\sum_{i=1}^{n}\gamma_{i}(\mathbf{X}_{i}-\mathbf{M}_{i})\right)-\lambda_{\max}\left(\sum_{i=1}^{n}\frac{\gamma_{i}^{2}}{6}(\mathbf{B}_{i}+2\mathbf{V}_{i})\right)\right\}.$ (74)
Proof.
Taking $\psi(x)=x^{2}/2$, $\mathbf{C}_{n}=\frac{1}{3}(\mathbf{X}_{n}-\mathbf{M})^{2}$, and $\mathbf{C}_{n}^{\prime}=\frac{2}{3}\mathbf{V}_{n}$ in Lemma 5.1, from (70),
$L_{n}^{\textsf{SN}}$ $\geqslant\exp\left\{\lambda_{\max}\left(\sum_{i=1}^{n}\gamma_{i}(\mathbf{X}_{i}-\mathbf{M}_{i})\right)-\lambda_{\max}\left(\sum_{i=1}^{n}\frac{\gamma_{i}^{2}}{6}((\mathbf{X}_{i}-\mathbf{M}_{i})^{2}+2\mathbf{V}_{i})\right)\right\}$ (75)
$\geqslant\exp\left\{\lambda_{\max}\left(\sum_{i=1}^{n}\gamma_{i}(\mathbf{X}_{i}-\mathbf{M}_{i})\right)-\lambda_{\max}\left(\sum_{i=1}^{n}\frac{\gamma_{i}^{2}}{6}(\mathbf{B}_{i}+2\mathbf{V}_{i})\right)\right\},$ (76)
concluding the proof. ∎
A closed-form, randomized, stopping time-valid matrix Bernstein or Hoeffding inequality on the concentration of weighted sample average is immediately in force by applying (UVI) to $\{L_{n}^{\textsf{SN}}\}$. Below, we denote by $\overline{\mathbf{X}}_{n}^{\gamma}$ the weighted average $\frac{\gamma_{1}\mathbf{X}_{1}+\dots+\gamma_{n}\mathbf{X}_{n}}{\gamma_{1}+\dots+\gamma_{n}}$ w.r.t. some positive weight sequence $\{\gamma_{n}\}$.
Theorem 5.5 (Uniformly Randomized Stopped Matrix Bernstein/Hoeffding Inequality).
Let $\{\mathbf{X}_{n}\}$ be $\mathcal{S}_{d}$-valued random matrices adapted to $\{\mathcal{F}_{n}\}$ with conditional means $\mathbb{E}(\mathbf{X}_{n}|\mathcal{F}_{n-1})=\mathbf{M}_{n}$, $(\mathbf{X}_{n}-\mathbf{M}_{n})^{2}\preceq\mathbf{B}_{n}$ and $\mathbb{E}((\mathbf{X}_{n}-\mathbf{M}_{n})^{2}|\mathcal{F}_{n-1})\preceq\mathbf{V}_{n}$ almost surely for predictable $\{\mathbf{B}_{n}\}$ and $\{\mathbf{V}_{n}\}$. Then, for any stopping time $\tau$, predictable positive sequence $\{\gamma_{n}\}$, $\alpha\in(0,1)$, and super-uniform random scalar $U$ independent of $\mathcal{F}_{\infty}$,
$\mathbb{P}\left(\lambda_{\max}\left(\overline{\mathbf{X}}_{\tau}^{\gamma}-\overline{\mathbf{M}}_{\tau}^{\gamma}\right)\geqslant\frac{\log(dU/\alpha)+\frac{1}{6}\lambda_{\max}\left(\sum{\gamma_{i}^{2}}(\mathbf{B}_{i}+2\mathbf{V}_{i})\right)}{\gamma_{1}+\dots+\gamma_{\tau}}\right)$$\leqslant$ (77)
$\mathbb{P}\left(\lambda_{\max}\left(\overline{\mathbf{X}}_{\tau}^{\gamma}-\overline{\mathbf{M}}_{\tau}^{\gamma}\right)\geqslant\frac{\log(dU/\alpha)+\frac{1}{2}\lambda_{\max}\left(\sum_{i=1}^{\tau}\gamma_{i}^{2}\mathbf{B}_{i}\right)}{\gamma_{1}+\dots+\gamma_{\tau}}\right)$ $\leqslant\alpha.$ (78)
We now remark at length on this theorem. First, the Hoeffding term $\frac{1}{2}\lambda_{\max}\left(\sum\gamma_{i}^{2}\mathbf{B}_{i}\right)$ comes from applying $\mathbf{V}_{i}\preceq\mathbf{B}_{i}$ to the Bernstein term $\frac{1}{6}\lambda_{\max}\left(\sum\gamma_{i}^{2}(\mathbf{B}_{i}^{2}+2\mathbf{V}_{i})\right)$, among which one may choose, as is with scalar Bernstein versus Hoeffding, depending on whether variance bounds $\{\mathbf{V}_{n}\}$ better than the total square bounds $\{\mathbf{B}_{n}\}$ are known. Second, the strength of this bound can be seen from the following fact: taking $\{\mathbf{M}_{n}\}$ to be a constant $\mathbf{M}$, $\{\mathbf{B}_{n}\}$ a constant $\mathbf{B}$, $\{\mathbf{V}_{n}\}$ a constant $\mathbf{V}$, $\tau$ a deterministic sample size $n$ and letting $U=1$, $\gamma_{1}=\dots=\gamma_{n}=\sqrt{\frac{6\log(d/\alpha)}{n\lambda_{\max}(\mathbf{B}+2\mathbf{V})}}$, (77) reads
$\mathbb{P}\left(\lambda_{\max}\left(\overline{\mathbf{X}}_{n}-\mathbf{M}\right)\geqslant\sqrt{\frac{2\log(d/\alpha)\lambda_{\max}(\mathbf{B}+2\mathbf{V})}{3n}}\right)\leqslant\alpha,$ (79)
where $\overline{\mathbf{X}}_{n}=\frac{1}{n}(\mathbf{X}_{1}+\dots+\mathbf{X}_{n})$. This even improves the optimal matrix Hoeffding inequality by Mackey et al. [2014, Corollary 4.2] whose dependence on $\mathbf{B}$ and $\mathbf{V}$ is $\frac{1}{2}\lambda_{\max}(\mathbf{B}+\mathbf{V})$ compared to our smaller $\frac{1}{3}\lambda_{\max}(\mathbf{B}+2\mathbf{V})$, [^4] The optimality of Mackey et al. [2014, Corollary 4.2] is in terms of the overall multiplicative constant which (79) achieves as well, not the distribution of the constant between $\mathbf{B}$ and $\mathbf{V}$. Even tighter dependence in this regard shall be presented soon in Section 5.3. and recovers (MHI), both via $\mathbf{V}\preceq\mathbf{B}$. The uncompromised inequality (77) further allows for stopping time-validity, time-varying and adaptive weighting, and randomized tightening simultaneously. Third, in the i.i.d. (or homoscedastic more generally) sequential (sample size $\tau$ not fixed) case, following Waudby-Smith and Ramdas [2020], Wang and Ramdas [2023a], the weights $\{\gamma_{n}\}$ can be set as a vanishing sequence such that each $\gamma_{n}$ matches the fixed-time optimal choice of $\gamma_{n}$ with sample size $n$, so we recommend $\gamma_{n}=\sqrt{\frac{6\log(d/\alpha)}{n\lambda_{\max}(\mathbf{B}+2\mathbf{V})}}$. By doing so, the concentration tightens itself as time goes on; or, the corresponding sequential test has asymptotic power 1 [Waudby-Smith and Ramdas, 2023, Table 1]. Moreover, it can be seen that the closed-form concentration of $\overline{\mathbf{X}}_{\tau}$ here is achieved via multiple steps of scalarization, each step loosening the bound by a little. Therefore, if one does not want to quantify uncertainty but is only interested in sequentially testing if the means $\mathbb{E}(\mathbf{X}_{n}|\mathcal{F}_{n-1})=\mathbf{M}$, the more powerful test in the current subsection is still with the largest process, the self-normalized supermartingale $\{L^{\textsf{SN}}_{n}\}$ (i.e., rejecting whenever $L^{\textsf{SN}}_{n}$ surpasses $Ud/\alpha$), with variance upper bounds $\mathbf{V}_{n}$ supplied by the upper bounds $\mathbf{B}_{n}$.
5.3 Matrix Empirical Bernstein Inequality
We now present a closed-form empirical Bernstein inequality that enjoys the same stopping time-validity, adaptive weighting, and randomization as well, foreshadowed by a matrix expectation bound in Proposition 4.10. Empirical Bernstein inequalities for scalars differ from “classical” Bernstein inequalities, e.g. our (77) when $d=1$, in that the population variance is replaced by an empirical variance in the deviation bound. They therefore achieve the advantage of Bernstein’s inequalities over Hoeffding’s inequalities without requiring knowing the variance beforehand.
Scalar empirical Bernstein inequalities are often derived in two different ways. First, a union bound between a classical Bernstein inequality and a concentration inequality on the empirical variance, a method that early empirical Bernstein results [Audibert et al., 2009, Maurer and Pontil, 2009] all share. Second, a self-normalization technique pioneered by Howard et al. [2021, Theorem 4] and later followed by Waudby-Smith and Ramdas [2023], Chugg et al. [2023], among others, which enables supermartingale-based, stopping time-valid concentration, and variance proxy by predictable estimates other than the empirical variance. These two methods are inherently different and as argued convincingly by Howard et al. [2021, Appendix A.8]: the latter’s avoidance of the union bound produces a sharper concentration. No matrix empirical Bernstein results, especially those in the second fashion, are previously known.
Our matrix extension of Howard et al. [2021, Theorem 4] starts from the bound we realier prove as Proposition 4.10. Recall that $\psi_{E}(x)=-\log(1-x)-x=\sum_{k=2}^{\infty}\frac{x^{k}}{k}$ for $x\in[0,1)$, and that
$\mathbb{E}\left(\exp(\gamma_{n}(\mathbf{X}_{n}-\widehat{\mathbf{X}}_{n})-\psi_{E}(\gamma_{n})(\mathbf{X}_{n}-\widehat{\mathbf{X}}_{n})^{2})\middle|\mathcal{F}_{n-1}\right)\preceq\exp(\gamma_{n}(\mathbf{M}_{n}-\widehat{\mathbf{X}}_{n}))$ (80)
under, among other, a key assumption that $\lambda_{\min}(\mathbf{X}_{n}-\widehat{\mathbf{X}}_{n})\geqslant-1$. We can use Lemma 5.1 again to produce a scalar supermartingale and a closed-form spectral concentration inequality.
Theorem 5.6 (Uniformly Randomized Stopped Matrix Empirical Bernstein Inequality).
Let $\{\mathbf{X}_{n}\}$ be $\mathcal{S}_{d}$-valued random matrices adapted to $\{\mathcal{F}_{n}\}$ with conditional means $\mathbb{E}(\mathbf{X}_{n}|\mathcal{F}_{n-1})=\mathbf{M}_{n}$. Let $\{\widehat{\mathbf{X}}_{n}\}$ be a sequence of predictable and integrable $\mathcal{S}_{d}$-valued random matrices such that $\lambda_{\min}(\mathbf{X}_{n}-\widehat{\mathbf{X}}_{n})\geqslant-1$ almost surely. Then, for any predictable $(0,1)$-valued sequence $\{\gamma_{n}\}$,
$L_{n}^{\mathsf{EB}}=\mathrm{tr}\exp\left(\sum_{i=1}^{n}\gamma_{i}(\mathbf{X}_{i}-\mathbf{M}_{i})-\sum_{i=1}^{n}\psi_{E}(\gamma_{i})(\mathbf{X}_{i}-\widehat{\mathbf{X}}_{i})^{2}\right)$ (81)
is a supermartingale. Further, for any stopping time $\tau$, $\alpha\in(0,1)$, and super-uniform scalar random variable $U$ independent of $\mathcal{F}_{\infty}$, we have
$\mathbb{P}\left(\lambda_{\max}\left(\overline{\mathbf{X}}_{\tau}^{\gamma}-\overline{\mathbf{M}}_{\tau}^{\gamma}\right)\geqslant\frac{\log(dU/\alpha)+\lambda_{\max}\left(\sum_{i=1}^{\tau}\psi_{E}(\gamma_{i})(\mathbf{X}_{i}-\widehat{\mathbf{X}}_{i})^{2}\right)}{\gamma_{1}+\dots+\gamma_{\tau}}\right)\leqslant\alpha.$ (82)
Before we remark on the uncompromised Theorem 5.6, let us first write down its fixed-time, non-randomized, fine-tuned special case with $\tau=n$ and $U=1$ which shall justify the “empirical Bernstein” name it bears.
Corollary 5.7 (Matrix Empirical Bernstein Inequality).
Suppose $\alpha\in(0,1)$. Let $\mathbf{X}_{1},\dots,\mathbf{X}_{n}$ be $\mathcal{S}_{d}^{[0,1]}$-valued i.i.d. random matrices with mean $\mathbf{M}=\mathbb{E}\mathbf{X}_{1}$. Let $\overline{\mathbf{X}}_{i}=\frac{1}{i}(\mathbf{X}_{1}+\dots+\mathbf{X}_{i})$ and $\overline{\mathbf{X}}_{0}=0$. Then, for any predictable $(0,1)$-valued sequence $\gamma_{1},\dots,\gamma_{n}$,
[…]
Further, if we define the following variance proxies
$\overline{\mathbf{V}}_{0}=0,\quad\overline{\mathbf{V}}_{k}=\frac{1}{k}\sum_{i=1}^{k}(\mathbf{X}_{i}-\overline{\mathbf{X}}_{k})^{2},\quad\overline{v}_{k}=\lambda_{\max}(\overline{\mathbf{V}}_{k})\vee\frac{5\log(d/\alpha)}{n},$ (84)
$\widetilde{s}_{n}=\lambda_{\max}\left(\frac{1}{n}\sum_{i=1}^{n}\frac{(\mathbf{X}_{i}-\overline{\mathbf{X}}_{i-1}^{2})}{\overline{v}_{i-1}}\right),$ (85)
and set $\gamma_{i}=\sqrt{\frac{2\log(d/\alpha)}{n\overline{v}_{i-1}}}$ for $i=1,\dots,n$, then,
$D_{n}^{\mathsf{EB}}\leqslant\sqrt{\frac{\log(d/\alpha)}{2n}}\frac{1+2\widetilde{s}_{n}}{\frac{1}{n}\sum_{i=1}^{n}\overline{v}_{i-1}^{-1/2}}\quad\approx\sqrt{\frac{9\log(d/\alpha)\lambda_{\max}(\mathbf{V})}{2n}}$ (86)
where $\mathbf{V}=\mathbf{Var}(\mathbf{X}_{1})$; further,
$\lim_{n\to\infty}\sqrt{n}D_{n}^{\mathsf{EB}}=\sqrt{2\log(d/\alpha)\lambda_{\max}(\mathbf{V})}\quad\text{almost surely}.$ (87)
We prove both Theorem 5.6 and Corollary 5.7 in Section B.4. The asymptotic behavior (87) of deviation bound $D_{n}^{\mathsf{EB}}$ is striking as it adapts fully to, without knowing, the true variance $\mathbf{V}$, a strict improvement even compared to the deviation bound $D_{n}^{\mathsf{B}}$ of the our previous matrix Bernstein inequality (79),
$\sqrt{n}D_{n}^{\mathsf{B}}=\sqrt{2\log(d/\alpha){\frac{1+2\lambda_{\max}(\mathbf{V})}{3}}},$ (88)
which requires knowing beforehand, but only adapts partially to, the variance $\mathbf{V}$. Indeed, when $\mathbf{V}$ is small, (88) can be arbitrarily larger than (87). We can also compare Corollary 5.7 to another matrix Bernstein inequality in the literature by Tropp [2012, Theorem 1.4], who proves that $\mathbb{P}\left(\lambda_{\max}\left(\overline{\mathbf{X}}_{n}-\mathbf{M}\right)\geqslant D_{n}^{\mathsf{TB}}\right)\leqslant\alpha$ under the same assumptions (plus knowing the variance $\mathbf{V}$), where its deviation bound
$D_{n}^{\mathsf{TB}}=\frac{\log(d/\alpha)+\sqrt{\log^{2}(d/\alpha)+18\log(d/\alpha)n\lambda_{\max}(\mathbf{V})}}{3n}$ (89)
satisfies the same asymptotic $\lim_{n\to\infty}\sqrt{n}D_{n}^{\mathsf{TB}}=\sqrt{2\log(d/\alpha)\lambda_{\max}(\mathbf{V})}$ as (87), which we achieve without requiring knowing $\mathbf{V}$. Other matrix Bernstein bounds in the literature, like the ones by Mackey et al. [2014, Corollary 5.2] and by Oliveira [2009], are looser than the Tropp [2012, Theorem 1.4] bound we discuss here as acknowledged by Mackey et al. [2014] themselves.
Our Theorem 5.6 and Corollary 5.7 owe much to the techniques developed by Waudby-Smith and Ramdas [2023, Theorem 2 and Remark 1] in the scalar case (who in turn build on the earlier result by Howard et al. [2021, Theorem 4] via the “predictable mixing” sequence $\{\gamma_{n}\}$). We recover exactly as special cases those two known results when $d=1$.
We reiterate our remark following Proposition 4.10 that $\widehat{\mathbf{X}}_{n}$ in Theorem 5.6 can usually be taken as the empirical average from $\mathbf{X}_{1}$ to $\mathbf{X}_{n-1}$, which is exactly what is done in Corollary 5.7. In the i.i.d. (or homoscedastic more generally) sequential case without the sample size $n$ fixed in advance, we can again (recalling a similar discussion following Theorem 5.5) set the weights $\{\gamma_{n}\}$ to be a vanishing sequence such that each $\gamma_{n}$ matches the fixed-time optimal choice of $\gamma_{n}$ with sample size $n$, so we recommend $\gamma_{n}=\sqrt{\frac{2\log(d/\alpha)}{n\overline{v}_{n-1}}}$. Under this weight sequence, we see the choice of a weighted average $\widehat{\mathbf{X}}_{n}=\overline{\mathbf{X}}_{n-1}^{\psi_{E}(\gamma)}$ is more reasonable as it roughly minimizes the weighted sum of squares $\sum_{i=1}^{\tau}\psi_{E}(\gamma_{i})(\mathbf{X}_{i}-\widehat{\mathbf{X}}_{i})^{2}$ in (82) into a weighted empirical variance — a heuristic which should work for Proposition 4.10 as well. Obviously, as long as $\widehat{\mathbf{X}}_{n}$ is any average, weighted or not, of $\mathbf{X}_{1},\dots,\mathbf{X}_{n-1}$, the condition $\lambda_{\min}(\mathbf{X}_{n}-\widehat{\mathbf{X}}_{n})\geqslant-1$ is met when $\{\mathbf{X}_{n}\}$ all take values in $\mathcal{S}_{d}^{[0,1]}$.
6 Experiments: To Scalarize or not to Scalarize
We have developed thus far two categories of matrix concentration bounds: those obtained via the pure matrix approach, typically expressed in the $\preceq$ order, in Sections 3 and 4; and those obtained via scalarization, i.e. taking the trace, in Section 5. It is therefore a natural question to ask if either of these is tighter than the other.
A fair comparison can be raised between the self-normalized matrix and scalar test supermartingales, Example 4.8 and Example 5.2 — for both are under the assumption of finite variance and based on the key one-step self-normalized bound (71). We place them in the setting of the following sequential testing problem: let $\mathbf{X}_{1},\mathbf{X}_{2},\dots$ be an i.i.d. sequence of matrices in $\mathcal{S}_{d}$, with mean $\mathbf{M}$ and variance upper bound $\mathbf{I}$. We would like to sequentially test the null hypothesis $H_{0}:\mathbf{M}=0$. Let us call the tests derived from Example 4.8 and Example 5.2 for this null Matrix and Scalar respectively. Ignoring the randomizer for simplicity (plug in $\mathbf{U}=\mathbf{I}$), Matrix computes the matrix-valued process
$\mathbf{Y}_{n}^{\textsf{SN}}=\exp\left(-\frac{\gamma_{1}^{2}+\dots+\gamma_{n}^{2}}{3}\right)\cdot\exp\left(\frac{\gamma_{1}\mathbf{X}_{1}}{2}-\frac{\gamma_{1}^{2}\mathbf{X}_{n}^{2}}{12}\right)\dots\exp\left(\frac{\gamma_{n}\mathbf{X}_{n}}{2}-\frac{\gamma_{n}^{2}\mathbf{X}_{n}^{2}}{12}\right)\\ \exp\left(\frac{\gamma_{n}\mathbf{X}_{n}}{2}-\frac{\gamma_{n}^{2}\mathbf{X}_{n}^{2}}{12}\right)\dots\exp\left(\frac{\gamma_{1}\mathbf{X}_{1}}{2}-\frac{\gamma_{1}^{2}\mathbf{X}_{1}^{2}}{12}\right),$ (90)
rejecting when $\mathbf{Y}_{n}^{\textsf{SN}}\npreceq\mathbf{A}$ for some threshold matrix $\mathbf{A}\in\mathcal{S}_{d}^{++}$ such that $\mathrm{tr}(\mathbf{A}^{-1})=\alpha$; Scalar computes the scalar-valued process
$L_{n}^{\textsf{SN}}=\exp\left(-\frac{\gamma_{1}^{2}+\dots+\gamma_{n}^{2}}{3}\right)\cdot\mathrm{tr}\exp\left(\sum_{i=1}^{n}\gamma_{i}\mathbf{X}_{i}-\sum_{i=1}^{n}\frac{\gamma_{i}^{2}\mathbf{X}_{i}^{2}}{6}\right),$ (91)
rejecting when $L_{n}^{\textsf{SN}}\geqslant d/\alpha$. Note that $L_{n}^{\textsf{SN}}$ does not equal $\mathrm{tr}(\mathbf{Y}_{n}^{\textsf{SN}})$ in general, an important point that we shall come to later.
Under the null, both Matrix and Scalar safely control the type 1 error (the probability of falsely rejecting the null) under the predefined level $\alpha$ due to (MVI) and (VI). For power comparison, we shall first discuss the case when only $n=1$ observation is revealed, and then move on to the case with an indefinite sequence of observations, with examples where Matrix outpowers Scalar under decreasingly restrictive conditions and finally no restriction at all.
6.1 Comparison with a Single Observation
When only $n=1$ observation is revealed, $L_{1}^{\textsf{SN}}=\mathrm{tr}(\mathbf{Y}_{1}^{\textsf{SN}})$. We shall see that the comparison depends heavily on the rejection threshold matrix $\mathbf{A}$ that Matrix picks beforehand.
Proposition 6.1.
Suppose Matrix picks $\mathbf{A}=(d/\alpha)\mathbf{I}$. Then, if Matrix rejects at $n=1$, Scalar rejects as well. In this case Scalar is at least as powerful as Matrix.
Proof.
If $\mathbf{A}=(d/\alpha)\mathbf{I}$ and Matrix rejects, then $\mathbf{Y}_{1}^{\textsf{SN}}\npreceq(d/\alpha)\mathbf{I}$, meaning that $\lambda_{\max}(\mathbf{Y}_{1}^{\textsf{SN}})>d/\alpha$. Therefore, $L_{1}^{\textsf{SN}}=\mathrm{tr}(\mathbf{Y}_{1}^{\textsf{SN}})\geqslant\lambda_{\max}(\mathbf{Y}_{1}^{\textsf{SN}})>d/\alpha$. ∎
Proposition 6.2.
Suppose that $d\geqslant 2$ and that the random matrix $\mathbf{Y}_{1}^{\mathsf{SN}}$ has full support, meaning it can equal any matrix in $\mathcal{S}_{d}^{++}$. Then,
  • $(d/\alpha)\mathbf{I}$ is the only matrix that Matrix can pick as $\mathbf{A}\in\mathcal{S}_{d}^{++}$ such that if Matrix rejects at $n=1$, Scalar always rejects as well. Therefore, as long as $\mathbf{A}\neq(d/\alpha)\mathbf{I}$, Scalar is not at least as powerful as Matrix;
  • It is impossible that if Scalar rejects at $n=1$, Matrix always rejects as well. So Matrix is never at least as powerful as Scalar.
In summary, for general $\mathbf{A}$ and $n=1$, neither Scalar nor Matrix dominates the other.
Proof.
For the first statement, consider the spectral decomposition of the matrix $\mathbf{A}\neq(d/\alpha)\mathbf{I}$,
$\mathbf{A}=\lambda_{1}\mathbf{P}_{1}+\dots+\lambda_{d}\mathbf{P}_{d},\quad\lambda_{1}\leqslant\dots\leqslant\lambda_{d}.$ (92)
The matrix $\lambda_{1}\mathbf{P}_{1}$ satisfies $\mathrm{tr}(\lambda_{1}\mathbf{P}_{1})=\lambda_{1}<d/\alpha$. Applying a small pertubation, $\lambda_{1}\mathbf{P}_{1}+\varepsilon\mathbf{I}\npreceq\mathbf{A}$, $\mathrm{tr}(\lambda_{1}\mathbf{P}_{1}+\varepsilon\mathbf{I})<d/\alpha$. So, if $\mathbf{Y}_{1}^{\mathsf{SN}}$ takes a value that is close to $\lambda_{1}\mathbf{P}_{1}+\varepsilon\mathbf{I}$, Matrix rejects at $n=1$ but Scalar does not.
For the second statement, the matrix $0.9\mathbf{A}$ satisfies $\mathrm{tr}(0.9\mathbf{A})\geqslant 0.9d^{2}/\alpha>d/\alpha$ by Cauchy-Schwarz inequality and $0.9\mathbf{A}\preceq\mathbf{A}$. Therefore, if $\mathbf{Y}_{1}^{\mathsf{SN}}$ takes a value that is close to $0.9\mathbf{A}$, Scalar rejects at $n=1$ but Matrix does not. ∎
Therefore, one may understand the difference between Matrix and Scalar as that Matrix has the freedom to choose the threshold matrix $\mathbf{A}$ in anticipation of the potential alternative; whereas Scalar represents some kind of Pareto optimal procedure if completely agnostic about the reality. For Matrix to be more powerful it has to choose $\mathbf{A}$ wisely according to some prior knowledge. The following example further showcases this point.
Example 6.3.
Suppose Matrix knows beforehand what $\mathbf{X}_{1}$ is (and hence what $\mathbf{Y}_{1}^{\mathsf{SN}}$ is), and $\lambda_{d}>1/\alpha$ where
$\mathbf{Y}_{1}^{\mathsf{SN}}=\lambda_{1}\mathbf{P}_{1}+\dots+\lambda_{d}\mathbf{P}_{d},\quad\lambda_{1}\leqslant\dots\leqslant\lambda_{d}$ (93)
is its spectral decomposition. Then, Matrix can always make sure to reject by picking
$\mathbf{A}=\varepsilon^{-1}(d-1)\mathbf{P}_{1}+\dots\varepsilon^{-1}(d-1)\mathbf{P}_{d-1}+(\alpha-\varepsilon)^{-1}\mathbf{P}_{d}$ (94)
with $\varepsilon\in(0,\alpha-\lambda_{d}^{-1})$. The condition $\lambda_{d}>1/\alpha$ here is strictly weaker than the condition under which Scalar rejects: $\lambda_{1}+\dots+\lambda_{d}=d/\alpha$.
Indeed, as we can see from these examples with $n=1$ (thus we have only compared the matrix Markov inequality with the scalar Markov inequality on the trace), for Matrix to outpower Scalar, apart from prior knowledge, it is favorable that $\lambda_{\max}(\mathbf{Y}_{1}^{\mathsf{SN}})$ is an outlier in the spectrum of $\mathbf{Y}_{1}^{\mathsf{SN}}$, i.e., ${\lambda_{\max}(\mathbf{Y}_{1}^{\mathsf{SN}})}/{\mathrm{tr}(\mathbf{Y}_{1}^{\mathsf{SN}})}$ is close to 1. This is exactly what we would predict from the discussion on (UMMI) early on, where we mention that (UMMI) is tighter when the random matrix has (approximately) a lower rank. In the very next subsection, we shall see that whereas the condition ${\lambda_{\max}(\mathbf{Y}_{n}^{\mathsf{SN}})}/{\mathrm{tr}(\mathbf{Y}_{n}^{\mathsf{SN}})}\approx 1$ seems artificial and contrived here with $n=1$, it becomes an almost certain phenomenon with large $n$.
6.2 Comparison with a Sequence of Observations
Let us consider first an example of a sequence of commuting observations, where Matrix’s excess power compared to Scalar increases as time goes on.
Example 6.4.
Suppose each observation has the spectral decomposition
$\mathbf{X}_{n}=\lambda_{1n}\mathbf{P}_{1}+\dots+\lambda_{dn}\mathbf{P}_{d}$ (95)
with common deterministic projection matrices $\mathbf{P}_{1},\dots,\mathbf{P}_{d}$. For each $k$, $\lambda_{k1},\lambda_{k2},\dots\overset{\mathrm{iid}}{\sim}\pi_{k}$ with mean $\mathbb{E}\lambda_{k1}=\lambda_{k}$ and variance $\mathbf{Var}\lambda_{k1}=\sigma_{k}^{2}\leqslant 1$. Then, the variance $\mathbf{Var}(\mathbf{X}_{1})=\sigma_{1}^{2}\mathbf{P}_{1}+\dots+\sigma_{d}^{2}\mathbf{P}_{d}\preceq\mathbf{I}$. The matrix-valued process that Matrix computes is
$\mathbf{Y}_{n}^{\textsf{SN}}=\sum_{k=1}^{d}Y_{n}^{\textsf{SN}}(\lambda_{k1},\dots,\lambda_{kn})\mathbf{P}_{k},$ (96)
and the scalar-valued process that Scalar computes is
$L_{n}^{\textsf{SN}}=\sum_{k=1}^{d}Y_{n}^{\textsf{SN}}(\lambda_{k1},\dots,\lambda_{kn})=\mathrm{tr}(\mathbf{Y}_{n}^{\textsf{SN}}).$ (97)
In both expressions the function $Y_{n}^{\textsf{SN}}(\cdot)$ denotes the self-normalized test process for scalar observations,
$Y_{n}^{\textsf{SN}}(x_{1},\dots,x_{n})=\exp\left(-\frac{1}{3}\sum_{i=1}^{n}\gamma_{i}^{2}+\sum_{i=1}^{n}\gamma_{i}x_{i}-\frac{1}{6}\sum_{i=1}^{n}\gamma_{i}^{2}x_{i}^{2}\right).$ (98)
The scalar component process $Y_{n}^{\textsf{SN}}(\lambda_{k1},\dots,\lambda_{kn})$ grows exponentially fast if $\pi_{k}$ has mean $\lambda_{k}\neq 0$, and shrinks to 0 if $\lambda_{k}=0$; moreover, the rate in which $Y_{n}^{\textsf{SN}}(\lambda_{k1},\dots,\lambda_{kn})$ grows depends exponentially on the size $|\lambda_{k}|$. Consequently, as long as there is any mean disparity between the component distributions $\pi_{1},\dots,\pi_{d}$, say $|\lambda_{1}|,\dots,|\lambda_{d-1}|<|\lambda_{d}|$, in the long run the process $\mathbf{Y}_{n}^{\textsf{SN}}$ will be approximately low-rank with ${\lambda_{\max}(\mathbf{Y}_{n}^{\mathsf{SN}})}/{\mathrm{tr}(\mathbf{Y}_{n}^{\mathsf{SN}})}\to 1$, as the component $Y_{n}^{\textsf{SN}}(\lambda_{d1},\dots,\lambda_{dn})\mathbf{P}_{d}$ will exponentially outgrow the rest. In this case if Matrix correctly anticipates the largest component $\mathbf{P}_{d}$, by picking again the threshold matrix
$\mathbf{A}=\varepsilon^{-1}(d-1)\mathbf{P}_{1}+\dots\varepsilon^{-1}(d-1)\mathbf{P}_{d-1}+(\alpha-\varepsilon)^{-1}\mathbf{P}_{d}$ (99)
with small $\varepsilon$ it will outperform Scalar: it only takes $Y_{n}^{\textsf{SN}}(\lambda_{d1},\dots,\lambda_{dn})$ to surpass $(\alpha-\varepsilon)^{-1}\approx 1/\alpha$ for Matrix to reject, whereas it takes $L_{n}^{\textsf{SN}}\approx Y_{n}^{\textsf{SN}}(\lambda_{d1},\dots,\lambda_{dn})$ to surpass $d/\alpha$ for Scalar to reject. Similar reasoning applies as long as Matrix knows a subset of the $d$ components with larger means.
We experiment with the following setup. We pick an orthonormal basis in $d=3$ dimension: $\mathbf{u}_{1}=(1/\sqrt{2},1/\sqrt{2},0)^{\mathsf{T}}$, $\mathbf{u}_{2}=(0,0,1)^{\mathsf{T}}$, $\mathbf{u}_{3}=(1/\sqrt{2},-1/\sqrt{2},0)^{\mathsf{T}}$, and let $\mathbf{P}_{1}=\mathbf{u}_{1}\mathbf{u}_{1}^{\mathsf{T}}$, $\mathbf{P}_{2}=\mathbf{u}_{2}\mathbf{u}_{2}^{\mathsf{T}}$, $\mathbf{P}_{3}=\mathbf{u}_{3}\mathbf{u}_{3}^{\mathsf{T}}$. The spectral distributions are set as $\pi_{1}=\mathcal{N}(0.1,1)$, $\pi_{2}=\mathcal{N}(-0.2,1)$, and $\pi_{3}=\mathcal{N}(0.3,1)$. We consider the following three sequential tests: smart Matrix, one that follows the recipe (99) with $\varepsilon=\alpha/20$; naive Matrix, one that is agnostic about the real distribution and picks $\mathbf{A}=(d/\alpha)\mathbf{I}$; and Scalar, all with the parameters $\gamma_{n}=n^{-1/2}$ and $\alpha=0.05$. We record the times at which these tests correctly reject the null, as well as the evolution of ${\lambda_{\max}(\mathbf{Y}_{n}^{\mathsf{SN}})}/{\mathrm{tr}(\mathbf{Y}_{n}^{\mathsf{SN}})}$. The findings are displayed in Figure 1: smart Matrix indeed outpowers the other two, whereas naive Matrix and Scalar are similar in power, explained by the quick convergence of ${\lambda_{\max}(\mathbf{Y}_{n}^{\mathsf{SN}})}/{\mathrm{tr}(\mathbf{Y}_{n}^{\mathsf{SN}})}$ to 1.
Figure 1: Empirical findings with commuting i.i.d. random matrices. Left. Comparison of the cumulative distribution functions of rejection times of the sequential tests corresponding to the matrix- and scalar-valued test processes under the same alternative distribution. Right. Evolution of the spectral distribution of the matrix-valued test process under the alternative.
While this indicates that
smart Matrix $>$ naive Matrix $\approx$ Scalar
in power, we now come to the striking (in light of Proposition 6.1) phenomenon that with the commutativity assumption removed, it is actually possible that
smart Matrix $>$ naive Matrix $>$ Scalar.
In the following example that involves non-commuting sequential observations, we consider the application scenario of covariance estimation [Ke et al., 2019].
Example 6.5.
Suppose $\mathbf{X}_{n}=\mathbf{x}_{n}\mathbf{x}_{n}^{\mathsf{T}}-(2d)^{-1}\mathbf{I}$ where $\mathbf{x}_{1},\mathbf{x}_{2},\dots\overset{\mathrm{iid}}{\sim}\mathcal{N}(0,\mathbf{\Sigma})$. We are to test if the null $\mathbf{\Sigma}=(2d)^{-1}\mathbf{I}$ is true, [^5] The null covariance $(2d)^{-1}\mathbf{I}$ is picked in the mere interest of simplicity so that the variance upper bound $\mathbf{Var}(\mathbf{X}_{n})\preceq\mathbf{I}$ is easily realizable in our example; it is easy to see that any $\mathbf{\Sigma}\in\mathcal{S}_{d}^{++}$ works as well by adjusting the variance upper bound accordingly. i.e., $\mathbb{E}\mathbf{X}_{n}=0$, under the 4th moment assumption $\mathbb{E}\|\mathbf{x}_{n}\|^{4}=(\mathrm{tr}\mathbf{\Sigma})^{2}+2\mathrm{tr}(\mathbf{\Sigma}^{2})\leqslant 1$, which implies
$\mathbf{Var}(\mathbf{X}_{n})=\mathbf{Var}(\mathbf{x}_{n}\mathbf{x}_{n}^{\mathsf{T}})\preceq\mathbb{E}(\mathbf{x}_{n}\mathbf{x}_{n}^{\mathsf{T}}\mathbf{x}_{n}\mathbf{x}_{n}^{\mathsf{T}})=\mathbb{E}(\|\mathbf{x}_{n}\|^{2}\cdot\mathbf{x}_{n}\mathbf{x}_{n}^{\mathsf{T}})\preceq\mathbb{E}(\|\mathbf{x}_{n}\|^{2}\cdot\|\mathbf{x}_{n}\|^{2}\mathbf{I})\preceq\mathbf{I}.$ (100)
Suppose in reality Matrix knows that $\mathbf{\Sigma}\neq(2d)^{-1}\mathbf{I}$ and $\mathbf{u}_{d}$ is an eigenvector of $\mathbf{\Sigma}-(2d)^{-1}\mathbf{I}$ with large eigenvalue, so that the data $\mathbf{X}_{1},\mathbf{X}_{2},\dots$ will manifest non-zero random signal approximately around this spectral component, it can first complete the orthonormal basis $\mathbf{u}_{1},\dots,\mathbf{u}_{d-1}$ and picks
$\mathbf{A}=\varepsilon^{-1}(d-1)\mathbf{u}_{1}\mathbf{u}_{1}^{\mathsf{T}}+\dots\varepsilon^{-1}(d-1)\mathbf{u}_{d-1}\mathbf{u}_{d-1}^{\mathsf{T}}+(\alpha-\varepsilon)^{-1}\mathbf{u}_{d}\mathbf{u}_{d}^{\mathsf{T}}$ (101)
to outpower Scalar.
Our experiment is set again in the $d=3$ case with $\mathbf{u}_{1}=(1/\sqrt{2},1/\sqrt{2},0)^{\mathsf{T}}$, $\mathbf{u}_{2}=(0,0,1)^{\mathsf{T}}$, $\mathbf{u}_{3}=(1/\sqrt{2},-1/\sqrt{2},0)^{\mathsf{T}}$, and the covariance matrix
$\mathbf{\Sigma}=\left(\begin{matrix}1/6&1/24&0\\ 1/24&1/6&0\\ 0&0&1/6\end{matrix}\right)=\text{null hypothesis}+\left(\begin{matrix}0&1/24&0\\ 1/24&0&0\\ 0&0&0\end{matrix}\right),$ (102)
which indeed satisfies $(\mathrm{tr}\mathbf{\Sigma})^{2}+2\mathrm{tr}(\mathbf{\Sigma}^{2})=\frac{61}{144}<1$. The three sequential tests as before are compared in Figure 2: the smart Matrix that follows (101) with $\varepsilon=\alpha/20$; the naive Matrix with $\mathbf{A}=(d/\alpha)\mathbf{I}$; Scalar, all with $\gamma_{n}=n^{-1/2}$ and $\alpha=0.05$. Besides the unsurprising fact that the smart Matrix is again the most powerful, rejecting the null earlier than the other two, we also observe the phenomenon that, while in the $n=1$ case, the matrix-valued test process with “naive” $\mathbf{A}=(d/\alpha)\mathbf{I}$ is always less powerful than the scalar-valued test process, this is reversed under non-commutativity with large $n$ as naive Matrix rejects slightly earlier than Scalar in our simulation. As also shown in Figure 2, the combination of two auxiliary findings contributes to this phenomenon: ${\lambda_{\max}(\mathbf{Y}_{n}^{\mathsf{SN}})}/{\mathrm{tr}(\mathbf{Y}_{n}^{\mathsf{SN}})}\to 1$ even when observations do not commute; and, while $L_{n}^{\textsf{SN}}=\mathrm{tr}(\mathbf{Y}_{n}^{\textsf{SN}})$ in all our previous examples, now $\mathrm{tr}(\mathbf{Y}_{n}^{\textsf{SN}})$ grows slightly faster than $L_{n}^{\textsf{SN}}$. It takes further theoretical investigations to understand why this happens (either in this particular example or in general), but to some extent it hints nonetheless at the possibility that in practice where the effect size is small (requiring large $n$) and matrix observations are non-commuting, there is an inherent advantage using the matrix-valued test process we propose.
Figure 2: Empirical findings with non-commuting i.i.d. random matrices for covariance testing. Left. Comparison of the cumulative distribution functions of rejection times of the sequential tests corresponding to the matrix- and scalar-valued test processes under the same alternative distribution. Middle. Evolution of the spectral distribution of the matrix-valued test process under the alternative. Right. Comparison between the trace of the matrix-valued test process, and the scalar-valued test process under the same alternative distribution.
7 Conclusion
We prove in this paper several novel concentration inequalities for random symmetric matrices. Some of these bounds are tighter than previous ones in the literature via randomization or Loewner order, others enjoy the virtues of validity at stopping times and/or the more general applicability to non-i.i.d. exchangeable matrices. These results shall find applications in, for example, covariance estimation and testing.
We also leave open some questions for future studies to address. These include a deeper understanding of the matrix randomizer $\mathbf{U}$ in our paper, which belongs to the “trace super-uniform” distribution family; tighter $p$-Chebyshev inequalities; further investigation into matrix-valued e-processes; as well as the theoretical explanation of our comparison between matrix-valued test supermartingales and scalar-valued test supermartingales when testing sequentially the same null hypotheses. Our experiments show that matrix-valued test processes are often more powerful in these sequential tests.
Acknowledgement
The authors thank Joel A. Tropp, Lester Mackey, and Tudor Manole for suggestions and discussions, and acknowledge support from NSF grant DMS-2310718.
Appendix A $p$-Chebyshev Inequalities
A.1 The Scalar (and Vector) $p$-Chebyshev Inequality
The following statement shows that for i.i.d. random scalars, the central $p$th moment of the sum grows linearly in the number of scalars just like the variance.
Proposition A.1 (Contraction of the scalar $p$th moment, Lemma 4.2 of Cherapanamjeri et al. [2022]).
Suppose $1\leqslant p\leqslant 2$. Let $X_{1},\dots,X_{n}$ be i.i.d. scalar with mean $\mu=\mathbb{E}X_{1}$ and central $p$th moment $v_{p}=\mathbb{E}|X_{1}-\mu|^{p}$. Then, $\mathbb{E}|X_{1}+\dots+X_{n}-n\mu|^{p}\leqslant 2^{2-p}nv_{p}$.
Wang et al. [2021] generalize the statement to vectors under the martingale difference assumption.
Proposition A.2 (Concentration of the vector $p$th moment, Lemma 7 of Wang et al. [2021]).
Suppose $1\leqslant p\leqslant 2$. Let $\mathbf{x}_{1},\dots,\mathbf{x}_{n}$ be $d$-dimensional random vectors adapted to a filtration $\mathcal{F}_{0},\mathcal{F}_{1},\dots,\mathcal{F}_{n}$ with conditional mean $\mu=\mathbb{E}(\mathbf{x}_{i}|\mathcal{F}_{i-1})$ and conditional central $p$th moment upper bound $v_{p}\geqslant\mathbb{E}(\|\mathbf{x}_{i}-\mu\|^{p}|\mathcal{F}_{i-1})$. Then, $\mathbb{E}\|\mathbf{x}_{1}+\dots+\mathbf{x}_{n}-n\mu\|^{p}\leqslant 2^{2-p}d^{1-p/2}nv_{p}$.
[…]
Corollary A.3 (Vector $p$-Chebyshev Inequality).
Suppose $\mathbf{x}_{1},\dots,\mathbf{x}_{n}$ satisfy the same assumptions as in Proposition A.2. Then, for any $a>0$,
$\mathbb{P}\left(\left\|\overline{\mathbf{x}}_{n}-\mu\right\|\geqslant a\right)\leqslant\frac{2^{2-p}d^{1-p}v_{p}}{n^{p-1}a^{p}}.$ (103)
A.2 Matrix $p$-Chebyshev Inequalities
The following conjecture is an ideal extension of Propositions A.2 and A.3.
Conjecture A.4 (Matrix $p$-Chebyshev Inequality).
Let $\mathbf{X}_{1},\dots,\mathbf{X}_{n}$ be i.i.d. random matrices taking values in $\mathcal{S}_{d}$ with common mean matrix $\mathbb{E}\mathbf{X}_{1}=\mathbf{M}$ and $p$th central moment matrix $\mathbf{V}_{p}:=\mathbb{E}(\operatorname{abs}(\mathbf{X}_{1}-\mathbf{M}))^{p}$. Then, for any $\mathbf{A}\in\mathcal{S}_{d}^{++}$, there exists a function $f:[1,2]\times\mathbb{N}\to(0,\infty)$ that grows sublinearly in its second argument, and bounded in its first argument, such that the $p$th central moment contracts,
$\mathbb{E}(\operatorname{abs}(\mathbf{X}_{1}+\dots+\mathbf{X}_{n}-n\mathbf{M})^{p})\preceq n\cdot f(p,d)\cdot\mathbf{V}_{p};$ (104)
and consequently, for $\overline{\mathbf{X}}_{n}=\frac{1}{n}(\mathbf{X}_{1}+\dots+\mathbf{X}_{n})$,
$\mathbb{P}(\operatorname{abs}(\overline{\mathbf{X}}_{n}-\mathbf{M})\npreceq\mathbf{A})\leqslant n^{1-p}\cdot f(p,d)\cdot\mathrm{tr}(\mathbf{V}_{p}\mathbf{A}^{-p}).$ (105)
If it holds, the $\mathbf{A}$ above can be randomized into $(\mathbf{A}^{p/2}\mathbf{U}\mathbf{A}^{p/2})^{1/p}$ as well.
We do have the following heavily scalarized and dimension-dependent bound that scales correctly w.r.t. the sample size $n$. We find this bound trivial as it regards the matrix as a mere collection of $d^{2}$ scalars and is strictly looser than the “standard” vector bound (Proposition A.2) applied to the flattened matrix.
Proposition A.5 (Matrix $p$-Chebyshev Spectral Inequality).
Let $\mathbf{X}_{1},\dots,\mathbf{X}_{n}$ be i.i.d. random matrices taking values in $\mathcal{S}_{d}$ with common mean matrix $\mathbb{E}\mathbf{X}_{1}=\mathbf{M}$ and $p$th spectral central moment $v_{p}:=\mathbb{E}\|\mathbf{X}_{1}-\mathbf{M}\|^{p}$. Then,
$\mathbb{E}\|\mathbf{X}_{1}+\dots+\mathbf{X}_{n}-n\mathbf{M}\|^{p}\leqslant 2^{2-p}d^{2-p/2}\cdot n\cdot v_{p}.$ (106)
Proof.
Without loss of generality suppose $\mathbf{M}=0$. Denote by $\|\cdot\|_{L_{p}}$ the entrywise matrix $p$-norm and $\|\cdot\|_{\mathrm{F}}=\|\cdot\|_{L_{2}}$ the Frobenius norm. Recall that $\|\mathbf{A}\|\leqslant\|\mathbf{A}\|_{\mathrm{F}}\leqslant\sqrt{d}\|\mathbf{A}\|$ and $\|\mathbf{A}\|_{\mathrm{F}}\leqslant\|\mathbf{A}\|_{L_{p}}\leqslant d^{\frac{2}{p}-1}\|\mathbf{A}\|_{\mathrm{F}}$. Then, $\mathbb{E}\|\mathbf{X}_{1}+\dots+\mathbf{X}_{n}\|^{p}$ is upper bounded by
$\mathbb{E}\|\mathbf{X}_{1}+\dots+\mathbf{X}_{n}\|_{L_{p}}^{p}=\mathbb{E}\left[\sum_{1\leqslant j,k\leqslant d}\left|\sum_{i=1}^{n}X_{i(jk)}\right|^{p}\right]=\sum_{1\leqslant j,k\leqslant d}\mathbb{E}\left|\sum_{i=1}^{n}X_{i(jk)}\right|^{p}$ (107)
$\stackrel{{\scriptstyle(*)}}{{\leqslant}}\,$ $\sum_{1\leqslant j,k\leqslant d}2^{2-p}n\mathbb{E}\left|X_{1(jk)}\right|^{p}=2^{2-p}n\mathbb{E}\|\mathbf{X}_{1}\|_{L_{p}}^{p}\leqslant 2^{2-p}d^{2-p}\cdot n\cdot\mathbb{E}\|\mathbf{X}_{1}\|^{p}_{\mathrm{F}}$ (108)
[…]
where the inequality (*)(*)( * ) comes from the contraction of scalar $p$th moment (Proposition A.1). ∎
Appendix B Additional Lemmas and Omitted Proofs
B.1 Proofs of Statements in Table 1
Proof.
For Rademacher, the bound $\mathbf{G}(\gamma)\preceq\exp(\gamma^{2}\mathbf{C}^{2}/2)$ is proved by Tropp [2015, Lemma 4.6.3] via a direct Taylor expansion argument. Since the matrix log is monotone (Fact 2), $\log\mathbf{G}(\gamma/n)\preceq\gamma^{2}\mathbf{C}^{2}/2n^{2}$. Finally, invoking the monotonicity of $\mathbf{A}\mapsto\mathrm{tr}\exp\mathbf{A}$ (Fact 3) we see that $\mathrm{tr}(\mathrm{e}^{n\log\mathbf{G}(\gamma/n)})\leqslant\mathrm{tr}(\mathrm{e}^{\gamma^{2}\mathbf{C}^{2}/2n})$. Bennett I follows from a similar reasoning applied to the Bennett MGF bound [Tropp, 2012, Lemma 6.7]. For Bennett II, we can apply the variance bound at last: $\mathrm{tr}(\mathrm{e}^{n(\mathrm{e}^{\gamma/n}-\gamma/n-1)\mathbf{Var}(\mathbf{X}_{1})})\leqslant\mathrm{tr}(\mathrm{e}^{n(\mathrm{e}^{\gamma/n}-\gamma/n-1)\mathbf{V}})$. The uni-Gaussian MGF bound is also proved by Tropp [2015, Lemma 4.6.2] via a direct Taylor expansion argument. For symmetric Hoeffding, introducing an independent $R\sim\operatorname{rad}_{1/2}$,
$\log\mathbf{G}(\gamma)=\log\mathbb{E}\mathrm{e}^{\gamma(\mathbf{X}_{1}-\mathbf{M})}=\log\mathbb{E}\mathrm{e}^{\gamma R(\mathbf{X}_{1}-\mathbf{M})}\preceq\gamma^{2}\mathbf{B}/2$ (110)
due to [Tropp, 2012, Lemma 7.7], leading to the corresponding upper bound on $\mathrm{tr}(\mathrm{e}^{\gamma^{2}\mathbf{C}^{2}/2n})$ as before. ∎
B.2 Dominated Integrability Lemma
It is well-known that if $X$ and $Y$ are scalar random variables such that $c\leqslant X\leqslant Y$ almost surely for some constant $c$ and that $\mathbb{E}|Y|<\infty$, it follows that $\mathbb{E}|X|<\infty$ as well, and $\mathbb{E}X\leqslant\mathbb{E}Y$. This type of “implied integrability” appears frequently in scalar concentration bounds. Let us prove its symmetric matrix extension for the sake of self-containedness.
Lemma B.1 (Dominated Integrability).
Let $\mathbf{X}$ and $\mathbf{Y}$ be $\mathcal{S}_{d}^{[c,\infty)}$-valued random matrices for some $c\in\mathbb{R}$ such that $\mathbf{X}\preceq\mathbf{Y}$ almost surely. Further, suppose $\mathbb{E}\mathbf{Y}$ exists. Then, so does $\mathbb{E}\mathbf{X}$ and $\mathbb{E}\mathbf{X}\preceq\mathbb{E}\mathbf{Y}$.
Proof.
Let us prove that each element $X_{ij}$ of the random matrix $\mathbf{X}$ is integrable. Note that for any deterministic $\mathbf{v}\in\mathbb{R}^{d}$, $\mathbf{v}^{\mathsf{T}}\mathbf{X}\mathbf{v}\leqslant\mathbf{v}^{\mathsf{T}}\mathbf{Y}\mathbf{v}$ almost surely. First, taking $\mathbf{v}=(0,\dots,0,1,0,\dots 0)^{\mathsf{T}}$, we have
$c\leqslant X_{jj}\leqslant Y_{jj}\quad\text{almost surely},$ (111)
concluding that the diagonal element $X_{jj}$ must be integrable (since $Y_{jj}$ is). Next, taking $\mathbf{v}=(0,\dots,0,1,0,\dots,0,1,0,\dots,0)^{\mathsf{T}}$, we have
$2c\leqslant 2X_{ij}+X_{ii}+X_{jj}\leqslant 2Y_{ij}+Y_{ii}+Y_{jj}\quad\text{almost surely},$ (112)
concluding that $2X_{ij}+X_{ii}+X_{jj}$ must be integrable (since $2Y_{ij}+Y_{ii}+Y_{jj}$ is). Therefore, the off-diagonal element $X_{ij}$ is integrable since $X_{ii}$ and $X_{jj}$ are.
Now that we have established the existence of $\mathbb{E}\mathbf{X}$, it is clear that $\mathbb{E}\mathbf{X}\preceq\mathbb{E}\mathbf{Y}$ since for any $\mathbf{v}\in\mathbb{R}^{d}$, $\mathbf{v}^{\mathsf{T}}(\mathbb{E}\mathbf{X})\mathbf{v}=\mathbb{E}(\mathbf{v}^{\mathsf{T}}\mathbf{X}\mathbf{v})\leqslant\mathbb{E}(\mathbf{v}^{\mathsf{T}}\mathbf{Y}\mathbf{v})=\mathbf{v}^{\mathsf{T}}(\mathbb{E}\mathbf{Y})\mathbf{v}$. ∎
B.3 Proofs of Self-Normalization Statements
The following transfer rule [Tropp, 2012, Equation 2.2] shall commonly be used: suppose $I\subseteq\mathbb{R}$ and $f,g:I\to\mathbb{R}$ satisfies $f(x)\leqslant g(x)$, then, since the function $g-f:I\to\mathbb{R}_{\geqslant 0}$ identifies canonically with a matrix-to-matrix function $g-f:\mathcal{S}_{d}^{I}\to\mathcal{S}_{d}^{+}$, it follows that $f(\mathbf{X})\preceq g(\mathbf{X})$ for any $\mathbf{X}\in\mathcal{S}_{d}^{I}$.
Proof of Example 4.8.
Let us replicate the key steps of the proof of Lemma 3(f) in Howard et al. [2020]. Recalling from Delyon [2009, Proposition 12] that $\exp(x-x^{2}/6)\leqslant 1+x+x^{2}/3$ for all $x\in\mathbb{R}$, we have, due to the transfer rule
$\exp(\gamma_{n}(\mathbf{X}_{n}-\mathbf{M}_{n})-\gamma_{n}^{2}(\mathbf{X}_{n}-\mathbf{M}_{n})^{2}/6)\preceq 1+\gamma_{n}(\mathbf{X}_{n}-\mathbf{M}_{n})+\gamma_{n}^{2}(\mathbf{X}_{n}-\mathbf{M}_{n})^{2}/3.$ (113)
This allows us to use Lemma B.1,
$\mathbb{E}\left(\exp(\gamma_{n}(\mathbf{X}_{n}-\mathbf{M}_{n})-\gamma_{n}^{2}(\mathbf{X}_{n}-\mathbf{M}_{n})^{2}/6)\mid\mathcal{F}_{n-1}\right)$ (114)
[…]
=\ = $1+\mathbb{E}\left(\gamma_{n}^{2}(\mathbf{X}_{n}-\mathbf{M}_{n})^{2}/3\mid\mathcal{F}_{n-1}\right)\preceq 1+\frac{\gamma_{n}^{2}}{3}\mathbf{V}_{n}\preceq\exp\left(\frac{\gamma_{n}^{2}}{3}\mathbf{V}_{n}\right),$ (116)
where in the last step we use the transfer rule again with $1+x\leqslant\exp(x)$ for all $x\in\mathbb{R}$. ∎
Proof of Example 4.9.
Let us replicate the key steps of the proof of Lemma 3(d) in Howard et al. [2020], which is also a self-normalization argument. Due to the symmetry of $\mathbf{X}_{n}-\mathbf{M}_{n}$, we introduce a Rademacher random variable $\epsilon$ independent from $\mathcal{F}_{n}$,
$\mathbb{E}\left(\exp\left(\gamma_{n}(\mathbf{X}_{n}-\mathbf{M}_{n})-\frac{\gamma_{n}^{2}(\mathbf{X}_{n}-\mathbf{M}_{n})^{2}}{2}\right)\middle|\mathcal{F}_{n-1}\right)$ (117)
$=$ $\mathbb{E}\left(\exp\left(\gamma_{n}\epsilon(\mathbf{X}_{n}-\mathbf{M}_{n})-\frac{\gamma_{n}^{2}(\mathbf{X}_{n}-\mathbf{M}_{n})^{2}}{2}\right)\middle|\mathcal{F}_{n-1}\right)$ (118)
$=$ $\mathbb{E}\left\{\mathbb{E}\left(\exp\left(\gamma_{n}\epsilon(\mathbf{X}_{n}-\mathbf{M}_{n})-\frac{\gamma_{n}^{2}(\mathbf{X}_{n}-\mathbf{M}_{n})^{2}}{2}\right)\middle|\mathcal{F}_{n}\right)\middle|\mathcal{F}_{n-1}\right\}$ (119)
$=$ $\mathbb{E}\left\{\frac{\exp\left(\gamma_{n}(\mathbf{X}_{n}-\mathbf{M}_{n})-\frac{\gamma_{n}^{2}(\mathbf{X}_{n}-\mathbf{M}_{n})^{2}}{2}\right)+\exp\left(-\gamma_{n}(\mathbf{X}_{n}-\mathbf{M}_{n})-\frac{\gamma_{n}^{2}(\mathbf{X}_{n}-\mathbf{M}_{n})^{2}}{2}\right)}{2}\middle|\mathcal{F}_{n-1}\right\}.$ (120)
Note that the function $f(x)=\frac{\exp(x-x^{2}/2)+\exp(-x-x^{2}/2)}{2}$ satisfies $f(x)\leqslant 1$ for all $x\in\mathbb{R}$, we see from the transfer rule and Lemma B.1 that
$\mathbb{E}\left(\exp\left(\gamma_{n}(\mathbf{X}_{n}-\mathbf{M}_{n})-\frac{\gamma_{n}^{2}(\mathbf{X}_{n}-\mathbf{M}_{n})^{2}}{2}\right)\middle|\mathcal{F}_{n-1}\right)=\mathbb{E}\left(f(\mathbf{X}_{n}-\mathbf{M}_{n})\middle|\mathcal{F}_{n-1}\right)\preceq\mathbf{I}.$ (121)
B.4 Proofs of Empirical Bernstein Statements
Proof of Proposition 4.10.
Recall that $\psi_{E}(\gamma)=-\log(1-\gamma)-\gamma$. An inequality by Fan et al. [2015] quoted by Howard et al. [2021, Appendix A.8] states that, for all $10\leqslant\gamma<$ 1 and $\xi\geqslant-1$,
$\exp(\gamma\xi-\psi_{E}(\gamma)\xi^{2})\leqslant 1+\gamma\xi.$ (122)
Since $\mathbf{X}_{n}-\widehat{\mathbf{X}}_{n}\in\mathcal{S}_{d}^{[-1,\infty)}$, we can apply the transfer rule mentioned in Section B.3, replacing the scalar $\xi$ above by the matrix $\mathbf{X}_{n}-\widehat{\mathbf{X}}_{n}$, and plugging in $\gamma=\gamma_{n}\in(0,1)$,
$\exp(\gamma_{n}(\mathbf{X}_{n}-\widehat{\mathbf{X}}_{n})-\psi_{E}(\gamma_{n})(\mathbf{X}_{n}-\widehat{\mathbf{X}}_{n})^{2})\preceq 1+\gamma_{n}(\mathbf{X}_{n}-\widehat{\mathbf{X}}_{n}).$ (123)
Lemma B.1 then guarantees the integrability of the left hand side, and that
$\mathbb{E}\left(\exp(\gamma_{n}(\mathbf{X}_{n}-\widehat{\mathbf{X}}_{n})-\psi_{E}(\gamma_{n})(\mathbf{X}_{n}-\widehat{\mathbf{X}}_{n})^{2})\middle|\mathcal{F}_{n-1}\right)\preceq\mathbb{E}\left(1+\gamma_{n}(\mathbf{X}_{n}-\widehat{\mathbf{X}}_{n})\middle|\mathcal{F}_{n-1}\right)$ (124)
$=$ $1+\gamma_{n}(\mathbf{M}_{n}-\widehat{\mathbf{X}}_{n})\preceq\exp(\gamma_{n}(\mathbf{M}_{n}-\widehat{\mathbf{X}}_{n})),$ (125)
where in the final step we use the transfer rule again with $1+x\leqslant\exp(x)$ for all $x\in\mathbb{R}$. This concludes the proof. ∎
Proof of Theorem 5.6.
We apply Lemma 5.1 with $\mathbf{Z}_{n}=\gamma_{n}(\mathbf{X}_{n}-\mathbf{M}_{n})$, $\mathbf{C}_{n}=\gamma_{n}(\widehat{\mathbf{X}}_{n}-\mathbf{M}_{n})+\psi_{E}(\gamma_{n})(\mathbf{X}_{n}-\widehat{\mathbf{X}}_{n})^{2}$, and $\mathbf{C}_{n}^{\prime}=\gamma_{n}(\mathbf{M}_{n}-\widehat{\mathbf{X}}_{n})$ to see that
$L_{n}^{\mathsf{EB}}=\mathrm{tr}\exp\left(\sum_{i=1}^{n}\gamma_{i}(\mathbf{X}_{i}-\mathbf{M}_{i})-\sum_{i=1}^{n}\psi_{E}(\gamma_{i})(\mathbf{X}_{n}-\widehat{\mathbf{X}}_{n})^{2}\right)$ (126)
is a supermartingale, which upper bounds
$\exp\left\{\lambda_{\max}\left(\sum_{i=1}^{n}\gamma_{i}(\mathbf{X}_{i}-\mathbf{M}_{i})\right)-\lambda_{\max}\left(\sum_{i=1}^{n}\psi_{E}(\gamma_{n})(\mathbf{X}_{n}-\widehat{\mathbf{X}}_{n})^{2}\right)\right\}.$ (127)
Applying (UVI),
$\mathbb{P}\left(\lambda_{\max}\left(\sum_{i=1}^{n}\gamma_{i}(\mathbf{X}_{i}-\mathbf{M}_{i})\right)-\lambda_{\max}\left(\sum_{i=1}^{n}\psi_{E}(\gamma_{n})(\mathbf{X}_{n}-\widehat{\mathbf{X}}_{n})^{2}\right)\geqslant\log(dU/\alpha)\right)\leqslant\alpha,$ (128)
and the desired result follows from rearranging. ∎
Lemma B.2.
Let $\psi_{E}(x)=-\log(1-x)-x$. When $0\leqslant x\leqslant\sqrt{2/5}$, $\psi_{E}(x)\leqslant x^{2}$.
Proof.
Let $g(x)=\psi_{E}(x)-x^{2}$. The claim follows from $g^{\prime\prime}(x)=(1-x)^{-2}-2\geqslant 0$ for $x\in[0,\sqrt{2/5}]$, and $g(0)=0$, $g(\sqrt{2/5})<0$. ∎
Proof of Corollary 5.7.
First, it is straightforward that $\lambda_{\max}({\mathbf{X}_{i}-\overline{\mathbf{X}}_{i-1}})\geqslant-1$ for every $i=1,\dots,n$ since both $\mathbf{X}_{i}$ and $\overline{\mathbf{X}}_{i-1}$ take values in $\mathcal{S}_{d}^{[0,1]}$, so Theorem 5.6 is applicable. Let us prove the two claims about the deviation bound $D_{n}^{\mathsf{EB}}$ under $\gamma_{i}=\sqrt{\frac{2\log(d/\alpha)}{n\overline{v}_{i-1}}}$. Recall that
$\overline{\mathbf{V}}_{0}=0,\quad\overline{\mathbf{V}}_{k}=\frac{1}{k}\sum_{i=1}^{k}(\mathbf{X}_{i}-\overline{\mathbf{X}}_{k})^{2},\quad\overline{v}_{k}=\lambda_{\max}(\overline{\mathbf{V}}_{k})\vee\frac{5\log(d/\alpha)}{n},$ (129)
$\widetilde{s}_{n}=\lambda_{\max}\left(\frac{1}{n}\sum_{i=1}^{n}\frac{(\mathbf{X}_{i}-\overline{\mathbf{X}}_{i-1}^{2})}{\overline{v}_{i-1}}\right).$ (130)
First, via our definition, $\overline{v}_{i}\geqslant\frac{5\log(d/\alpha)}{n}$, therefore $\gamma_{i}\leqslant\sqrt{2/5}$. So Lemma B.2 implies that $\psi_{E}(\gamma_{i})\leqslant\gamma_{i}^{2}$. Therefore,
$D_{n}^{\mathsf{EB}}$ $=\frac{\log(d/\alpha)+\lambda_{\max}\left(\sum_{i=1}^{n}\psi_{E}\left(\sqrt{\frac{2\log(d/\alpha)}{n\overline{v}_{i-1}}}\right)(\mathbf{X}_{i}-\overline{\mathbf{X}}_{i-1})^{2}\right)}{\sum_{i=1}^{n}\sqrt{\frac{2\log(d/\alpha)}{n\overline{v}_{i-1}}}}$ (131)
$\leqslant\frac{\log(d/\alpha)+\lambda_{\max}\left(\sum_{i=1}^{n}{\frac{2\log(d/\alpha)}{n\overline{v}_{i-1}}}(\mathbf{X}_{i}-\overline{\mathbf{X}}_{i-1})^{2}\right)}{\sum_{i=1}^{n}\sqrt{\frac{2\log(d/\alpha)}{n\overline{v}_{i-1}}}}$ (132)
$=\sqrt{\frac{\log(d/\alpha)}{2n}}\frac{1+2\widetilde{s}_{n}}{\frac{1}{n}\sum_{i=1}^{n}\overline{v}_{i-1}^{-1/2}}.$ (133)
Then, let us show that $\lim_{n\to\infty}D_{n}^{\mathsf{EB}}/D_{n}^{\mathsf{B}}$ almost surely. Via the boundedness of $\mathbf{X}_{1}$ and the strong law of large numbers, we can see the following limits hold almost surely (cf. Waudby-Smith and Ramdas [2023, Lemmas 4-6]):
$\lim_{k\to\infty}\overline{\mathbf{X}}_{k}=\mathbf{M},\quad\lim_{k\to\infty}\overline{\mathbf{V}}_{k}=\mathbf{V}=\mathbf{Var}(\mathbf{X}_{1}),$ (134)
$\lim_{k\to\infty}\overline{v}_{k}=\lambda_{\max}(\mathbf{V}),\quad\lim_{n\to\infty}\frac{1}{n}\sum_{i=1}^{n}{(\mathbf{X}_{i}-\overline{\mathbf{X}}_{i-1})^{2}}=\mathbf{V},$ (135)
$\lim_{n\to\infty}\lambda_{\max}\left(\frac{1}{n}\sum_{i=1}^{n}{(\mathbf{X}_{i}-\overline{\mathbf{X}}_{i-1})^{2}}\right)=\lambda_{\max}(\mathbf{V}),$ (136)
$\lim_{n\to\infty}\lambda_{\max}\left(\frac{1}{n}\sum_{i=1}^{n}\frac{(\mathbf{X}_{i}-\overline{\mathbf{X}}_{i-1})^{2}}{\overline{v}_{i-1}}\right)=1.$ (137)
Therefore, we have, via the expansion $\psi_{E}(x)=\sum_{k=2}^{\infty}\frac{x^{k}}{k}$,
$\limsup_{n\to\infty}\sqrt{n}D_{n}^{\mathsf{EB}}$ (138)
$=$ $\limsup_{n\to\infty}\frac{\log(d/\alpha)+\lambda_{\max}\left(\sum_{i=1}^{n}\psi_{E}\left(\sqrt{\frac{2\log(d/\alpha)}{n\overline{v}_{i-1}}}\right)(\mathbf{X}_{i}-\overline{\mathbf{X}}_{i-1})^{2}\right)}{\frac{1}{n}\sum_{i=1}^{n}\sqrt{\frac{2\log(d/\alpha)}{\overline{v}_{i-1}}}}$ (139)
$\leqslant$ $\limsup_{n\to\infty}\frac{\log(d/\alpha)+\lambda_{\max}\left(\sum_{i=1}^{n}\frac{1}{2}\left(\sqrt{\frac{2\log(d/\alpha)}{n\overline{v}_{i-1}}}\right)^{2}(\mathbf{X}_{i}-\overline{\mathbf{X}}_{i-1})^{2}\right)}{\frac{1}{n}\sum_{i=1}^{n}\sqrt{\frac{2\log(d/\alpha)}{\overline{v}_{i-1}}}}$ (140)
$+$ $\sum_{k=3}^{\infty}\limsup_{n\to\infty}\frac{\lambda_{\max}\left(\sum_{i=1}^{n}\frac{1}{k}\left(\sqrt{\frac{2\log(d/\alpha)}{n\overline{v}_{i-1}}}\right)^{k}(\mathbf{X}_{i}-\overline{\mathbf{X}}_{i-1})^{2}\right)}{\frac{1}{n}\sum_{i=1}^{n}\sqrt{\frac{2\log(d/\alpha)}{\overline{v}_{i-1}}}}$ (141)
$=$ $\limsup_{n\to\infty}\frac{\log(d/\alpha)+\lambda_{\max}\left(\sum_{i=1}^{n}\frac{1}{2}\left(\sqrt{\frac{2\log(d/\alpha)}{n\overline{v}_{i-1}}}\right)^{2}(\mathbf{X}_{i}-\overline{\mathbf{X}}_{i-1})^{2}\right)}{\frac{1}{n}\sum_{i=1}^{n}\sqrt{\frac{2\log(d/\alpha)}{\overline{v}_{i-1}}}}$ (142)
$=$ $\limsup_{n\to\infty}\sqrt{\frac{\log(d/\alpha)}{2}}\frac{\left(1+\lambda_{\max}\left(\frac{1}{n}\sum_{i=1}^{n}\frac{(\mathbf{X}_{i}-\overline{\mathbf{X}}_{i-1})^{2}}{\overline{v}_{i-1}}\right)\right)}{\frac{1}{n}\sum_{i=1}^{n}\overline{v}_{i-1}^{-1/2}}$ (143)
$=$ $\sqrt{2\log(d/\alpha)\lambda_{\max}(\mathbf{V})}.$ (144)
Similarly, one can show that $\liminf_{n\to\infty}\sqrt{n}D_{n}^{\mathsf{EB}}\geqslant\sqrt{2\log(d/\alpha)\lambda_{\max}(\mathbf{V})}$, concluding the proof. ∎
B.5 Proof of Lemma 5.1
Proof.
Due to the monotonicity of $\log$ (Fact 2), the condition (68) implies
$\log\mathbb{E}(\exp(\mathbf{Z}_{n}-\mathbf{C}_{n})|\mathcal{F}_{n-1})\preceq\mathbf{C}_{n}^{\prime}.$ (145)
Now recall Lieb’s concavity theorem [Lieb, 1973]: for any $\mathbf{H}\in\mathcal{S}_{d}$, the map $\mathbf{X}\mapsto\mathrm{tr}\exp(\mathbf{H}+\log\mathbf{X})$ ($\mathcal{S}_{d}^{++}\to(0,\infty)$) is concave. Therefore,
$\mathbb{E}(L_{n}|\mathcal{F}_{n-1})=$ $\mathbb{E}\left(\mathrm{tr}\exp\left(\sum_{i=1}^{n-1}\mathbf{Z}_{i}-\sum_{i=1}^{n-1}(\mathbf{C}_{i}+\mathbf{C}_{i}^{\prime})-\mathbf{C}_{n}^{\prime}+\log\mathrm{e}^{\mathbf{Z}_{n}-\mathbf{C}_{n}}\right)\middle|\mathcal{F}_{n-1}\right)$ (146)
(Jensen’s inequality)
[…]
(by (145) and Fact 3)
[…]
concluding the proof that $\{L_{n}\}$ is a supermartingale. Finally, observe that
$L_{n}$ $=\mathrm{tr}\exp\left(\sum_{i=1}^{n}\mathbf{Z}_{i}-\sum_{i=1}^{n}(\mathbf{C}_{i}+\mathbf{C}_{i}^{\prime})\right)$ (149)
$\geqslant\mathrm{tr}\exp\left(\sum_{i=1}^{n}\mathbf{Z}_{i}-\lambda_{\max}\left(\sum_{i=1}^{n}(\mathbf{C}_{i}+\mathbf{C}_{i}^{\prime})\right)\mathbf{I}\right)$ (150)
$\geqslant\lambda_{\max}\exp\left(\sum_{i=1}^{n}\mathbf{Z}_{i}-\lambda_{\max}\left(\sum_{i=1}^{n}(\mathbf{C}_{i}+\mathbf{C}_{i}^{\prime})\right)\mathbf{I}\right)$ (151)
$=\exp\lambda_{\max}\left(\sum_{i=1}^{n}\mathbf{Z}_{i}-\lambda_{\max}\left(\sum_{i=1}^{n}(\mathbf{C}_{i}+\mathbf{C}_{i}^{\prime})\right)\mathbf{I}\right)$ (152)
$=\exp\left(\lambda_{\max}\left(\sum_{i=1}^{n}\mathbf{Z}_{i}\right)-\lambda_{\max}\left(\sum_{i=1}^{n}(\mathbf{C}_{i}+\mathbf{C}_{i}^{\prime})\right)\right),$ (153)
concluding the proof. ∎
Appendix C Further Notes on Matrix-Valued Martingales
C.1 Doob’s $L^{p}$ Inequality on Trace
Applying an operator-monotone and concave function to a matrix supermartingale, we get another matrix supermartingale. In particular,
Lemma C.1.
Let $\{\mathbf{M}_{n}\}$ be an $\mathcal{S}_{d}^{+}$-valued supermartingale on $\{\mathcal{F}_{n}\}$ and $p\in[1,\infty)$. Then, $\{\mathbf{M}_{n}^{1/p}\}$ and $\{\log\mathbf{M}_{n}\}$ are also supermartingales.
Proof.
The function $\mathbf{X}\mapsto\mathbf{X}^{1/p}$ is operator-monotone and concave. So
$\mathbb{E}(\mathbf{M}_{n}^{1/p}|\mathcal{F}_{n-1})\preceq\mathbb{E}(\mathbf{M}_{n}|\mathcal{F}_{n-1})^{1/p}\preceq\mathbf{M}_{n-1}^{1/p}.$ (154)
The same proof applies for the logarithm as well. ∎
Fact 3 and Fact 4, on the other hand, convert a matrix-valued (super/sub)martingale to a scalar-valued one. For example,
Lemma C.2.
Let $\{\mathbf{M}_{n}\}$ be an $\mathcal{S}_{d}^{+}$-valued submartingale or $\mathcal{S}_{d}$-valued martingale, and $p\geqslant 1$. Then, $\{\mathrm{tr}(\operatorname{abs}\mathbf{M}_{n})^{p}\}$ is a $\mathbb{R}_{\geqslant 0}$-valued submartingale.
Proof.
Using Fact 4,
$\mathbb{E}(\mathrm{tr}(\operatorname{abs}\mathbf{M}_{n})^{p}|\mathcal{F}_{n-1})=\mathrm{tr}\mathbb{E}((\operatorname{abs}\mathbf{M}_{n})^{p}|\mathcal{F}_{n-1})\geqslant\mathrm{tr}(\operatorname{abs}\mathbb{E}(\mathbf{M}_{n}|\mathcal{F}_{n-1}))^{p}.$ (155)
If $\{\mathbf{M}_{n}\}$ is a martingale, the right hand side above is $\mathrm{tr}(\operatorname{abs}\mathbf{M}_{n-1})^{p}$. If $\mathbf{M}_{n}\succeq 0$ is a submartingale, then we have by Fact 3,
$\mathrm{tr}(\operatorname{abs}\mathbb{E}(\mathbf{M}_{n}|\mathcal{F}_{n-1}))^{p}\geqslant\mathrm{tr}(\operatorname{abs}\mathbf{M}_{n-1})^{p},$ (156)
since the function $x\mapsto|x|^{p}$ is monotone on $x\geqslant 0$. ∎
Proposition C.3 (Trace Doob’s $L^{p}$ Inequality).
Let $\{\mathbf{M}_{n}\}$ be an $\mathcal{S}_{d}^{+}$-valued submartingale or $\mathcal{S}_{d}$-valued martingale, and $p\geqslant 1$. Then,
$\mathbb{E}\left(\sup_{0\leqslant n\leqslant N}\|\mathbf{M}_{n}\|_{\mathrm{op}}^{p}\right)\leqslant\mathbb{E}\left(\sup_{0\leqslant n\leqslant N}(\mathrm{tr}\operatorname{abs}\mathbf{M}_{n})^{p}\right)\leqslant\left(\frac{p}{p-1}\right)^{p}\mathbb{E}(\mathrm{tr}\operatorname{abs}\mathbf{M}_{N})^{p}\leqslant\left(\frac{pd}{p-1}\right)^{p}\mathbb{E}\|\mathbf{M}_{N}\|_{\mathrm{op}}^{p}.$ (157)
Proof.
The second inequality is directly from applying Doob’s $L^{p}$ inequality on the nonnegative submartingale $\{\mathrm{tr}(\operatorname{abs}\mathbf{M}_{n})\}$. The first and the third $\leqslant$’s are both trivial. ∎
Appendix D Non-symmetric, Non-square Matrices
We finally remark on the extension of our methods to non-symmetric and possibly non-square matrices. Denote by $\mathcal{M}_{d_{1}\times d_{2}}(\mathbb{R})$ the space of all $d_{1}\times d_{2}$ real-valued matrices, and we consider the random observations $\mathbf{X}_{1},\mathbf{X}_{2},\dots$ now as taking values in $\mathcal{M}_{d_{1}\times d_{2}}(\mathbb{R})$. We can employ the classical idea of dilations (see e.g. [Tropp, 2015, Section 2.1.16]). Consider the linear map
$D:\mathcal{M}_{d_{1}\times d_{2}}(\mathbb{R})\to\mathcal{S}_{d_{1}+d_{2}},\quad D(\mathbf{X})=\left(\begin{matrix}0&\mathbf{X}\\ \mathbf{X}^{\mathsf{T}}&0\end{matrix}\right),$ (158)
and we can apply our previously developed bounds or tests on $\{D(\mathbf{X}_{n})\}$, which shall in turn provide concentration inequalities for $\{\mathbf{X}_{n}\}$ themselves: either via the relation $\|D(\mathbf{X})\|=\|\mathbf{X}\|$ with the linearity of $D$ to obtain spectral bounds for $\mathbf{X}_{1},\mathbf{X}_{2},\dots$, or one may directly use the resulting (scalar or matrix) e-process as the test process.
In Table 2, we see that quadratic assumptions on the possibly non-square random matrix $\mathbf{X}_{1}$ translate directly to those on the symmetric random matrix $D(\mathbf{X}_{1})$.
Assumption on $\mathbf{X}_{1}$ Assumption on $D(\mathbf{X}_{1})$
$\mathbf{X}_{1}^{\mathsf{T}}\mathbf{X}_{1}\preceq\mathbf{B}_{1}$ and $\mathbf{X}_{1}\mathbf{X}_{1}^{\mathsf{T}}\preceq\mathbf{B}_{2}$ $D^{2}(\mathbf{X}_{1})\preceq\left(\begin{matrix}\mathbf{B}_{1}&\\ &\mathbf{B}_{2}\end{matrix}\right)$
$\mathbf{Cov}(\mathbf{X}_{1}^{\mathsf{T}})\preceq\mathbf{V}_{1}$ and $\mathbf{Cov}(\mathbf{X}_{1})\preceq\mathbf{V}_{2}$ $\mathbf{Var}(D(\mathbf{X}_{1}))\preceq\left(\begin{matrix}\mathbf{V}_{1}&\\ &\mathbf{V}_{2}\end{matrix}\right)$
Table 2: Distribution assumptions on $\{\mathbf{X}_{n}\}$ imply those on $\{D(\mathbf{X}_{n})\}$. Here, the covariance of a random matrix is defined as $\mathbf{Cov}(\mathbf{X})=\mathbb{E}(\mathbf{X}-\mathbb{E}\mathbf{X})(\mathbf{X}-\mathbb{E}\mathbf{X})^{\mathsf{T}}$.
For example, the self-normalized matrix supermartingale (Example 4.8) can be generalized as follows.
Proposition D.1.
Let $\{\mathbf{X}_{n}\}$ be $\mathcal{M}_{d_{1}\times d_{2}}(\mathbb{R})$-valued random matrices adapted to $\{\mathcal{F}_{n}\}$ with conditional means $\mathbb{E}(\mathbf{X}_{n}|\mathcal{F}_{n-1})=\mathbf{M}_{n}$ and conditional covariance upper bounds $\mathbf{Cov}(\mathbf{X}_{n}^{\mathsf{T}}|\mathcal{F}_{n-1})\preceq\mathbf{V}_{1n}$, $\mathbf{Cov}(\mathbf{X}_{n}|\mathcal{F}_{n-1})\preceq\mathbf{V}_{2n}$ that are predictable. For any predictable positive scalars $\{\gamma_{n}\}$, let
$\scriptsize\mathbf{E}_{n}=\left(\begin{matrix}-\gamma_{n}^{2}(\mathbf{X}_{n}-\mathbf{M}_{n})(\mathbf{X}_{n}-\mathbf{M}_{n})^{\mathsf{T}}/6&\gamma_{n}(\mathbf{X}_{n}-\mathbf{M}_{n})\\ \gamma_{n}(\mathbf{X}_{n}-\mathbf{M}_{n})^{\mathsf{T}}&-\gamma_{n}^{2}(\mathbf{X}_{n}-\mathbf{M}_{n})^{\mathsf{T}}(\mathbf{X}_{n}-\mathbf{M}_{n})/6\end{matrix}\right),\;\mathbf{F}_{n}=-\frac{\gamma_{n}^{2}}{3}\left(\begin{matrix}\mathbf{V}_{1n}&\\ &\mathbf{V}_{2n}\end{matrix}\right),$ (159)
the process
$\mathbf{Y}_{n}=\mathrm{e}^{\mathbf{F}_{1}/2}\mathrm{e}^{\mathbf{E}_{1}/2}\mathrm{e}^{\mathbf{F}_{2}/2}\mathrm{e}^{\mathbf{E}_{2}/2}\dots\mathrm{e}^{\mathbf{F}_{n}/2}\mathrm{e}^{\mathbf{E}_{n}/2}\mathrm{e}^{\mathbf{E}_{n}/2}\mathrm{e}^{\mathbf{F}_{n}/2}\dots\mathrm{e}^{\mathbf{E}_{2}/2}\mathrm{e}^{\mathbf{F}_{2}/2}\mathrm{e}^{\mathbf{E}_{1}/2}\mathrm{e}^{\mathbf{F}_{1}/2}.$ (160)
is a $\mathcal{S}_{d_{1}+d_{2}}^{+}$-valued supermartingale.